### v1

# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

Abstract
Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images. However, existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities. In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1,162 images across hundreds of object categories and attack words. Through extensive benchmarking of Vision-Language Models (VLMs) on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks. Our findings reveal that typographic attacks persist in state-of-the-art Large Vision-Language Models (LVLMs) due to the choice of their vision encoder, though larger Large Language Models (LLMs) backbones help mitigate their vulnerability. Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research. Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems. We publicly release the datasets introduced in this paper under https://huggingface.co/datasets/BLISS-e-V/SCAM, along with the code for evaluations at https://github.com/Bliss-e-V/SCAM.
Figure 1: Overview of the SCAM datasets and benchmarking framework. A) We create three counterfactual versions of images: SCAM, NoSCAM, and SynthSCAM. B) Evaluation of VLMs via cosine similarity between image and text. C) Prompt-based evaluation of LVLMs.
1 Introduction
VLMs such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. With the growing use of VLMs in real-world applications, such as medical applications [10, 11, 12] and autonomous systems [13], as well as in Large Vision-Language Models (LVLMs) [14, 15, 16], it is essential to ensure the safety and reliability of these models. However, recent research has revealed that these models are vulnerable to typographic attacks [17, 18, 19], which exploit the reliance on the textual component of multimodal learning.
Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions. Figure 1 depicts how adding the word “taxi” to an image of a “clock” can cause a model to misclassify it as a “taxi”. These attacks expose a fundamental interaction in multimodal learning: the model gets tricked by focusing on text within an image, more than the visual content itself [17, 18, 19]. Understanding and mitigating such vulnerabilities is essential for the safe deployment of multimodal foundation models.
To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest real-world typographic attack dataset to date. Each image contains an object with a semantically unrelated handwritten attack word [^1] While we use the term “word”, our attacks also incorporate phrases like “no left turn”. on a post-it note. Furthermore, we provide two variations, a clean dataset where attack words are removed (NoSCAM), and a synthetic dataset where attack words are digitally reintroduced (SynthSCAM).
We provide extensive evaluations over a range of VLMs and LVLMs and show that:
  1. 1.
    Models of both classes experience a significant drop in accuracy when typographic attacks are introduced (SCAM).
  2. 2.
    We find that synthetic attacks closely align with real-world attacks, providing empirical support for the established practice of evaluating on synthetic datasets.
  3. 3.
    LVLMs, such as the LLaVA family, inherit typographic vulnerabilities from their vision encoders.
  4. 4.
    Utilizing larger Large Language Model (LLM) backbones can alleviate an LVLM’s inherited weaknesses regarding typographic robustness.
Table 1: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM datasets. Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding. Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [20]. This table presents a selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
Model Training data Accuracy (%)
NoSCAM SCAM
RN50 openai 97.50 36.32 ↓61.18
ViT-B-32 laion2b 97.68 74.18 ↓23.50
ViT-B-16 laion2b 98.54 68.93 ↓29.61
ViT-B-16-SigLIP webli 99.05 81.50 ↓17.55
ViT-L-14 commonpool_xl 99.40 74.96 ↓24.44
openai 98.80 41.05 ↓57.75
ViT-L-14-336 openai 99.23 34.68 ↓64.55
ViT-g-14 laion2b 98.71 62.05 ↓36.66
ViT-bigG-14 laion2b 99.40 70.91 ↓28.49
llava-llama3-1.1:8b - 97.83 40.17 ↓57.66
llava-1.6:7b - 96.30 57.78 ↓38.52
llava-1.6:13b - 97.86 62.42 ↓35.44
llava-1.6:34b - 98.62 86.67 ↓11.95
llama3.2-vision:90b - 99.05 70.75 ↓28.30
gpt-4o-mini-2024-07-18 - 99.14 91.39 ↓7.75
gpt-4o-2024-08-06 - 99.57 97.41 ↓2.16
2 Related Work
Multimodal Foundation Models.
Multimodal pretraining, as established by CLIP [1] and SigLIP [2], leverages image-text pairs to capture rich semantics suitable for a variety of applications [21, 22]. This approach commonly features a contrastive objective to align visual and textual content in a shared latent space. Following this paradigm, various Vision-Language Models (VLMs) [23, 24, 25] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7]. Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [26, 27, 28, 29, 28]. Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
Typographic Attacks.
While vision-language pretraining enables strong performance, research has shown that VLMs and LVLMs are susceptible to typographic attacks [17, 18, 19]. Existing works demonstrate several approaches to employ typographic attacks, for instance by Visual-Prompt-Injection [30], with self-generated attacks [19], or in scene-coherent fashion [31]. Since VLMs are integrated into autonomous systems, typographic attacks are also known to occur in safety-critical applications such as autonomous driving [32].
Datasets.
To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works [33, 34, 18, 19, 30, 35] synthetically add attack words into existing image datasets while only few [36, 37, 38] propose datasets that exhibit real-world typographic attacks. Existing real-world datasets, such as PAINT [38] and Materzynska+[37], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models. The previously largest real-world typographic attack dataset, RTA-100 [36], consists of 1,000 images spanning 100 objects and attack words. However, existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see table 2 for a comparison). Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains. This gap hinders not only the rigorous evaluation of VLM s and LVLM s but also the development of effective defense mechanisms.
Figure 2: Examples of images from the SCAM dataset. Additional images for both NoSCAM and SynthSCAM are provided in fig. 7 in the appendix.
Figure 3: Distribution of attack words in SCAM into categories, highlighting both everyday terms and safety-critical vocabulary.
3 SCAM Dataset
We introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date, containing 1,162 images (see table 2). The dataset and its two variations (NoSCAM and SynthSCAM), allowing for an empirical assessment of transferability between these settings, are publicly available on Hugging Face, along with the fully compatible and publicly accessible SCAM benchmark repository for standardized evaluation.
Data Collection.
We collected 1,162 images from diverse environments, such as households, utility stores, and traffic infrastructure, ensuring a varied and representative dataset. An overview of example images from SCAM is presented in fig. 2. Each image in the SCAM dataset captures an object together with an attack word – unrelated in meaning – handwritten on a single post-it note in English and placed alongside the object, as illustrated in fig. 1A.
To create the cleaned NoSCAM dataset, we remove all attack words by obscuring the post-it notes while preserving the original object images. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font) and matching the colors of the post-its and ink from the original images.
Data Annotation and Processing.
All images were manually annotated with both object and corresponding attack labels as part of a voluntary community effort. The annotation process was conducted using Label Studio [39]. To standardize the dataset, all images were zero-padded and downscaled to 512×512 pixels using cv2.INTER_AREA [40], an interpolation method that preserves details while minimizing distortion.
Dataset Statistics.
SCAM comprises 1,162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations. We provide lists of the 660 unique object labels and 206 attack words in section A.2. We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems. The distribution of these groups is depicted in fig. 3.
As shown in table 2, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words. Compared to RTA-100 [36], SCAM covers a much broader range of objects (660 vs. 100) and attack words (206 vs. 100), offering a more comprehensive evaluation of model robustness to typographic attacks. This increased diversity makes SCAM a stronger benchmark for assessing the vulnerabilities of multimodal foundation models.
Table 2: Comparison of real-world typographic attack datasets. SCAM is much more diverse in terms of objects and attack words.
Dataset Images Objects Attack words
PAINT [38] 110 89 87
Materzynska+ [37] 180 20 8
RTA-100 [36] 1,000 100 100
SCAM (ours) 1,162 660 206
4 Evaluation
We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup. Both evaluations are conducted on the SCAM, NoSCAM and SynthSCAM datasets. The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
Experimental Setup.
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [41, 42, 1, 43, 44]. We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 28, 29, 28] and Llama3.2-vision [45] models. The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API. An overview of selected models is provided in table 1, while LABEL:tab:full-overview presents the complete list of all 106 evaluated models.
Evaluating VLMs.
We evaluate the performance of VLMs in a zero-shot classification task on the SCAM datasets. For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word. These text embeddings are generated using the set of text templates proposed in [1], as detailed in section A.3. The predicted label is then determined based on the highest cosine similarity score. An overview of this process is depicted in fig. 1B. Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly. All evaluations of VLMs were performed on an A100 40GB card.
Evaluating LVLMs.
To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications. Our evaluation follows the setup illustrated in fig. 1C, inspired by [47]. The model is provided with an image from SCAM, NoSCAM, or SynthSCAM, along with the following prompt:
What entity is depicted in the image?
(a) clock
(b) taxi
Respond strictly with only (a) or (b), nothing else."
[…]
5 Results
We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table 1. Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points. LVLMs exhibit similar behavior and we find that the choice of vision encoder plays a critical role in determining the extent of accuracy degradation.
Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig. 4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table 3. These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
Figure 4: Accuracy distribution of $9$ VLMs across NoSCAM, SCAM, and SynthSCAM datasets. SCAM effectively lowers accuracy, and its similarity to SynthSCAM suggests synthetic attacks replicate real ones. LABEL:tab:full-overview in the appendix provides a full list of models and their accuracies.
Table 3: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks. The tables present average confusion matrices. The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks. Values represent the mean ± standard deviation, rounded to integers.
((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
[…]
((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
[…]
Zero-Shot Results for VLMs.
As shown in fig. 4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset. However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets. Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 35 percentage points. For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
Impact of the Training Dataset and Model Architecture
We investigate the vulnerability of VLMs by examining the impact of their training data, training paradigm, and model architecture, on their susceptibility to typographic attacks.
Table 4 presents the impact of training data on the ViT-B models [48] trained by the OpenCLIP team. We focus solely on these models as they share the same training paradigm but are trained on different datasets. Overall, the ViT-B models trained on the LAION datasets [41] are highly susceptible to typographic attacks, while certain CommonPool variants [49] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively. Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI. Consequently, as indicated by the accuracies for that model in table 1, this filtering criterion appears to propagate susceptibility.
Table 4: Training dataset choice significantly affects accuracy, as shown by the results on SCAM for the model ViT-B-16 trained by the OpenCLIP team on different datasets.
Dataset Accuracy %
commonpool_l_text 91.57
commonpool_l_image 90.71
commonpool_l_basic 89.76
commonpool_l 88.47
datacomp_l 79.60
commonpool_l_laion 78.83
commonpool_l_clip 70.91
laion2b 68.93
laion400m 68.80
datacomp_xl 68.07
Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to as v1). More details about the selection process can be in section A.5.
Training data Architecture Accuracy %
openai CLIP-ViT 44.06
CLIP-RN 35.92
v1 SigLIP-ViT 75.77
CLIP-ViT 69.02
While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance. We compare different architectures: CLIP-RN (ResNets [51]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [50] (see table 5). Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models. Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
Our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table 6. Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
Table 6: Models become more susceptible to typographic attacks as patch size decreases. Dataset fixed to laion2b, models evaluated on SCAM.
Model Patch Size Accuarcy %
ViT-B-32 32 74.18
ViT-bigG-14 14 70.91
ViT-B-16 16 68.93
ViT-H-14 14 63.08
ViT-L-14 14 62.91
ViT-g-14 14 62.05
Lastly, fig. 5 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks. However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
Figure 5: Susceptibility to typographic attack (evaluation on SCAM) is agnostic of VLMs size, measured in millions of parameters.
Impact of the Size of the Attack.
We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig. 6). The mean accuracy across all models follows this trend, which is consistent with the findings in [34].
Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean ± standard deviation across all 99 VLMs that we evaluate.
Prompt-Based Classification Results for LVLMs.
Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility. Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 30-50 percentage points under typographic attacks. This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig. 6). However, as model size increases, the robustness to typographic attacks improves. LLaVA models with larger LLM backbones, such as llava-1.6:34b, exhibit significantly better performance under attack. This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions. In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 2 percentage points. The smaller gpt-4o-mini, however, experiences a higher drop of approximately 8 points, illustrating again that larger models tend to be more robust. Despite that, model size alone does not determine robustness. In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava-1.6:34b, highlighting the critical role of the vision encoder in overall performance. These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
6 Discussion
Our evaluations on the Subtle Character Attacks on Multimodal Models (SCAM) datasets confirm previous findings [17, 18, 19] that multimodal foundation models remain vulnerable to typographic attacks, whether real-world (SCAM) or synthetic (SynthSCAM). Notably, we find that synthetic attacks are just as effective as real-world ones, leading to comparable performance drops.
Our analysis further reveals that training regime, training data, and model architecture significantly influence susceptibility of Vision-Language Models (VLMs) to typographic attacks, whereas model size alone does not correlate with robustness. In Large Vision-Language Models (LVLMs), vulnerability to typographic attacks carry over from the vision encoder, making these models inherently susceptible. However, we find that larger Large Language Model (LLM) backbones can mitigate this effect, leading to improved robustness and accuracy to typographic attacks.
Limitations.
A key limitation of our work is the use of “naive prompting” [52, 53] for LVLMs. Utilizing multi-prompt evaluation, such as those proposed in [54, 47], could provide a better assessment of their robustness to typographic attacks. Another limitation is the focus on English-language attacks and it should be assessed whether language-specific factors influence model susceptibility.
Future Work.
As multimodal models are increasingly deployed to safety-critical applications [12, 11, 13, 13], ensuring robustness against typographic attacks is essential to prevent serious consequences. Approaches such as alignment training [55] or defense techniques [36] offer promising directions for enhancing robustness. Additionally, our findings suggest that an LVLM ’s vulnerability to typographic attacks is closely tied to its vision encoder. To investigate this further, we plan to retrain LLaVA models with different vision backbones to systematically evaluate their impact on robustness. This will help determine whether stronger vision encoders mitigate susceptibility or if additional fine-tuning strategies are required.
7 Conclusion
In this paper, we introduced the Subtle Character Attacks on Multimodal Models (SCAM) dataset (and two variations, NoSCAM and SynthSCAM), the largest and most diverse collection of real-world images designed to evaluate the vulnerability of multimodal foundation models to typographic attacks. Our evaluation demonstrated that current state-of-the-art multimodal models are significantly affected by typographic attacks, highlighting a critical weakness in their robustness. We hope that our broad analysis of Vision-Language Models (VLMs) will help researchers make more informed choices when selecting typographically robust models for their downstream applications.
Acknowledgments and Disclosure of Funding
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12. We thank the BLISS [^2] https://bliss.berlin/ community – its very existence, collaborative energy, and shared commitment to open research made this work possible. Special thanks to Zeynep Altayli, Jonas Loos, Philippa Ringe, and Joseph Tschörner for their valuable efforts in image collection and labeling. We especially thank Jonas Loos for his significant work on the project page.
Author Contributions
JW was responsible for image processing, code development, model evaluations, visualizations, and overall coordination. EP contributed to conceptualization, figure design, and manuscript writing, including the introduction, dataset, evaluation, and drafted most of the results and discussion sections. JH focused on related work research and supported overall writing. LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback. LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work. All authors contributed to image collection and labeling.
Appendix A Appendix
A.1 SCAM Examples
In fig. 2, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
Figure 7: Example images of SCAM, NoSCAM, and SynthSCAM.
A.2 SCAM Labels
The SCAM dataset includes diverse object labels:
30kmh traffic zone sign, adapter, advent calendar, alarm clock, analog timer, apple, apple seed cutter, apples, armchair, ashtray, aubergine, avocado, axe, back, backpack, bacon bits, badminton ball, bag, baking pan, ball, balloon, banana, bananas, band-aid, barricade, baseball cap, basket, basket ball, bathroom sign, bathtub, batteries, bed, bell, bell pepper, bench, bench vise, bicycle, bicycle bell, bicycle light, bicycle tire, bike, bike light, bike lock, bin, bird feeder, blackboard, blanket, blender, blower, board, boat figurine, book, boot, bottle, bottle cap, bottle crate, bottle opener, bottles, boulder, bowl, box, boxes, brass tubes, bread slicer, bricks, broom, brooms, brush, brushes, bucket, butterfly, buttons, cabinet, cable, cable drum, cable ties, cage, cake, calculator, calendar, camera, candle, candles, candy cane, canister, cans, cap, capacitor, car, carafe, cardboard box, carrot, cash register, cashew nuts, cat figurine, cat tower, cauldron, celery, chain, chain saw, chair, chalk bag, champagne, charger, cheese, chess board, chess piece, chess set, chessboard, chewing gum, chewing gums, chili pepper, chips, chocolate, chopsticks, cigarettes, cleaner, clipboard, clock, closet, clothes iron, clothespin, coat, coat hanger, coat hook, coffee beans, coffee machine, coffee maker, coin, coin counter, combination lock, compass, condom, condoms, confetti, contact grill, controller, cooking scraper, cordless drill, cork, corn, counter, cow meat, crepe pan, crochet hook, cross wrench, cucumber, cup, deodorant, detergent, detergent powder, diapers, dice, diffuser, dish scrubber, dish soap, dishwasher, dishwasher tablets, disinfectant dispenser, dispenser, diving goggles, dog prohibition sign, doll, door, doorhandle, doormat, drill, drills, drink, drone, drum, duct tape, dustpan, dvd, e-scooter, ear phones, ear plugs, ear protection, easel, edamame, edding, eggs, electric kettle, electric screwdriver, electric shaver, electric toothbrush, electrical panel, elephant, emergency exit sign, end of parking zone sign, ernie, exit sign, faucet, feeder, fence, fennel, fever thermometer, figurine, fineliner pen, fire alarm button, fire detector, fire extinguisher, firework, first aid kit, fish, flip flops, flipper, floss, flower, flower pot, flowering pot, flowers, folder, folding rule, fork, fortune cookie, freeze pack, french press, fridge, fries, frisbee, frothing pitcher, fryer, funnel, garden gnome, garlic, gas bottle, gas cooker, gem lettuce, ginger, glas, glass, glass bottle, glasses, globe, glove, glue stick, gnocchi, goal, goggles, goose, grapes, grater, grave candles, green beans, grill, guitar, guitar foot stand, hair clip, hair clipper, hair dryer, hair iron, hairbrush, hairgrip, hairspray, ham, hammer, hand blender, hand brush and dustpan, handbrush, handkerchiefs, handwash, hanger, hard disk, hare, hat, hdmi splitter, headband, headphones, heater, helmet, high visibility jacket, hook, horn, hot dog maker, hot glue gun, hot water bottle, hotplate, ice cube tray, ice pack, ice scraper, icecream, intercom, interdental brush, iron, ironing board, jacket, jam, jar, jersey, juggling ball, juicer, kettle, key, key chain, keyboard, keys, kiwi, knife, labeling machine, ladder, lamp, landing net, lantern, laptop, lawn chair, letter box, lid, light bulbs, light switch, lightbulb, lighter, lime, lipstick, lollies, lollipop, loofah, macbook, magazine, makeup powder, mango, marker, masher, mask, matryoshka doll, meatball, melon, memory card, menstrual pad, menstrual pads, mercedes-benz, metronome, microcontroller, microphone, microwave, milk pitcher, mirror, moka pot, money, monitor, mosquito repellent, mosquito spray, motion detector, mouse, mug, mulled wine, multiple socket, mushroom, mushrooms, nail clipper, nail polish, napkin, necklace, no entry sign, no parking sign, no pedestrian crossing sign, no smoking sign, no-entry sign, noodle cooker, notebook, nuts, oat bar, oats, octopus, office chair, olives, onion, onions, orange, organizer, oven, oven glove, paint, paint brush, paint roller, painting, pair, palette knife, pallet, pan, panini press, pants, paper rolls, paper streamers, paper towel, paper towels, paper wrap, paprika, parking sign, parking ticket machine, parsley, party hat, pasta, pea, pen, pencil, pepper, pepper caster, perfume, person, phone, pick, pickaxe, piggy bank, pillow, pills, pine cone, pineapple plant, pineapples, ping pong ball, pipe connection, pipes, pizza, pizza cutter, pizza oven, pizza peel, plane ticket, plant, plastic cups, plate, playing card, plectrum, plug board, plug socket, plunger, pocket balance, pocket knife, pokemon ball, pole, portable electric drill, post, post-its, pot, potato, potato chips, potato masher, potatoes, pouch, pouffe, power drill, powerbank, printer, printer cartridge, projector, puffed rice, pump, pumpkin, puzzle, puzzle piece, rack, racket, radiator, radio, railing, rake, razor, record player, remote, remote control, resistor, ribbon, rice cooker, roll board, roll-on, roller, roller board, rolling pin, rope, rose, rosemary, roses, router, rubber bands, rubber duck, rubik’s cube, running shoes, safe, safety pin, safety pins, safety sign, sage leaves, salad, salami, salmon, salt, sandal, sandpaper, sandwich maker, santa figurine, sausage, sausages, saw, scissor, scissors, scooter, screen, screw, screwdriver, screws, seafood salad, serving tray, serving trolley, sewing machine, shaker, shampoo, shashlik skewers, shaver, shelf, shell, shirt, shoe, shoes, shopping cart, shovel, shovels, shower, shower gel, shower head, shrimp, shuttlecock, sieve, sign, silverware box, sink, skillet, skimmer, skimmers, sledgehammer, sleeping mat, slider, slipper, slippers, sneaker, soap, soap dispenser, sock, socket, socks, spade, speaker, spice mill, spinach, spiral notebooks, sponge, spoon, spout, spray bottle, spray can, spring, squeegee, stamp, stapler, steam flat iron, step bin, stereo, stone plates, stool, stop sign, storage box, straw, straws, string, stuffed animal, suitcase, sunglasses, sunscreen, sweeper, t-shirt, table, takeaway bag, tampon, tap, tape, tape dispenser, tea, tea pot, teabag, tealight, tee box, telephone, tennis ball, terminal, thermal camping mat, thermometer, thread, ticket machine, tin box, tire, tissues, toaster, toilet, toilet bowl, toilet brush, toilet lid, toilet paper, tomato, tomatoes, tongue brush, toothbrush, toothpaste, tortellini, towel, toy car, toy elephant, toy horse, traffic light, traffic signs, trailer, trash bag, trash bin, trash can, trashcan, tray, tree, try square, turmeric, turtle, tv, ufonauts, umbrella, underpants, usb cable, usb-c cable, vacuum, vacuum cleaner, vase, vegetable peeler, video surveillance sign, vinyl, waffles, wall, wall charger, wallet, washbasin, washing machine, watch, water wings, watering can, watermelon, webcam, weight, wet floor sign, wheel, whisk, white cabbage, wildlife camera, window, window cleaner, wine, wood, wooden disc, wooden grill tong, wooden panels, wooden pestle, wooden skewer, wool, wrapping paper, wrench, yarn, and zucchini.
Table 7 summarizes the frequency of object labels. The top 15 labels (e.g., “lamp”, “plant”, “bottle”) are among the most frequent, while the bottom 5 labels (e.g., “chain”, “roller board”, “mask”) appear rarely. Notably, 437 object labels appear exactly once, illustrating a pronounced long-tail distribution that underscores the diversity of the dataset.
Table 7: Top 15 and bottom 5 object labels by appearance frequency.
Object label Count
lamp 13
plant 13
bottle 13
chair 11
glass 10
candle 10
cable 8
grill 8
printer 8
cup 8
speaker 8
pillow 8
brush 7
canister 7
pen 7
chain 1
roller board 1
mask 1
goggles 1
webcam 1
SCAM includes a variety of attack words:
apple, arrow, bag, ball, barrier, basketball, beach, beatles, bed, beer, bell, bench, bicycle, bike, bird, blanket, boat, book, boot, bottle, boulder, bowl, box, boy, bracelet, bridge, bridge out, broom, bus, butter, cable, cake, camera, canal, candle, cane, cannon, cap, car, carpet, carrot, cart, cat, cellphone, chair, cheese, chest, chicken, child, clock, cloud, coat, coin, cone, couch, cow, crab, crystal, cup, curb, curtaxin, curve, dead end, deer, desert, desk, detour, dish, dog, door, duck, eagle, egg, elevator, exit, fan, feather, fence, finger, fire, fish, flag, floor, flower, fog, forest, fork, frog, fruit, garbage bin, gears, ginger, gloves, grandpa, grape, grass, gravel, green, grenade, gun, hair, hammer, hat, helmet, highway, hill, hole, honey, horn, horse, house, ice, iron, jacket, jar, jeep, key, knife, ladder, lake, lamp, lane, leaf, lemon, lentils, lift, lightbulb, lion, lipstick, lock, macbook, mad, median, merge, mirror, mittens, mountaxin, no left turn, no parking, ocean, olive oil, oregano, pedestrian, pen, person, phone, piano, picasso, pig, police, pumpkin, raccoon, red light, risotto rice, river, road, road closed, rock, saddle, salt, sandwich, santa claus, scarf, scooter, screw, screwdriver, sea, seed, shadow, ship, sidewalk, signal, slippery, sloth, slow, smartphone, sofa, speed camera, speed limit, spok, spoon, staxirs, stop, stop sign, store, street, street sign, suv, table, taxi, toll, traffic light, traffic sign, train, transistor, trash, tree, truck, unicorn, van, vase, wall, wet, woman, yield, and zebra.
Table 8 summarizes the frequency of attack words. The top 15 attack words (e.g., “knife”, “truck”, “slow”) occur frequently, whereas many attack words are rare. In fact, 56 attack words appear exactly once, again highlighting overall diversity of the dataset.
Table 8: Top 15 and bottom 5 attack words by appearance frequency.
Attack Word Count
knife 24
truck 23
slow 20
lion 20
zebra 19
helmet 18
bike 17
honey 15
pig 15
camera 15
lemon 15
pen 15
tree 14
house 14
fish 13
beatles 1
green 1
raccoon 1
beach 1
unicorn 1
A.3 Zero-Shot Evaluation of LVMs: Text Templates
To compute the cosine similarity between an image and text embeddings in our zero-shot evaluation of Vision-Language Models (VLMs), we generate text embeddings for both the object label and the attack word using multiple textual variations. Specifically, we apply 32 diverse text templates as in [1], each incorporating the object label resp. attack word in a different context. The final embedding is obtained by averaging the embeddings of all templates. The templates are as follows: ( ’a bad photo of a {}.’, ’a photo of many {}.’, ’a sculpture of a {}.’, ’a photo of the hard to see {}.’, ’a low resolution photo of the {}.’, ’a rendering of a {}.’, ’graffiti of a {}.’, ’a bad photo of the {}.’, ’a cropped photo of the {}.’, ’a tattoo of a {}.’, ’the embroidered {}.’, ’a photo of a hard to see {}.’, ’a bright photo of a {}.’, ’a photo of a clean {}.’, ’a photo of a dirty {}.’, ’a dark photo of the {}.’, ’a drawing of a {}.’, ’a photo of my {}.’, ’the plastic {}.’, ’a photo of the cool {}.’, ’a close-up photo of a {}.’, ’a black and white photo of the {}.’, ’a painting of the {}.’, ’a painting of a {}.’, ’a pixelated photo of the {}.’, ’a sculpture of the {}.’, ’a bright photo of the {}.’, ’a cropped photo of a {}.’, ’a plastic {}.’, ’a photo of the dirty {}.’, ’a jpeg corrupted photo of a {}.’, ’a blurry photo of the {}.’ ).
A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM datasets.
Model Training data Accuracy (%)
NoSCAM SCAM SynthSCAM
RN101 openai 97.42 40.02 27.71
RN50 openai 97.50 36.32 24.27
RN50x16 openai 98.62 37.69 35.28
RN50x4 openai 97.85 33.05 29.69
RN50x64 openai 99.05 32.53 32.70
ViT-B-16 commonpool_l_basic_s1b_b8k 98.54 89.76 82.87
commonpool_l_clip_s1b_b8k 98.11 70.91 63.08
commonpool_l_image_s1b_b8k 98.19 90.71 86.14
commonpool_l_laion_s1b_b8k 97.42 78.83 67.30
commonpool_l_s1b_b8k 97.50 88.47 81.33
commonpool_l_text_s1b_b8k 98.02 91.57 85.97
datacomp_l_s1b_b8k 98.02 79.60 69.62
datacomp_xl_s13b_b90k 98.80 68.07 55.34
laion2b_s34b_b88k 98.54 68.93 52.75
laion400m_e31 98.02 68.93 52.93
laion400m_e32 98.11 68.67 53.10
metaclip_400m 97.85 58.78 37.95
metaclip_fullcc 98.71 61.10 42.17
openai 97.07 40.02 22.03
ViT-B-16-SigLIP webli 99.05 81.50 76.25
ViT-B-16-SigLIP-256 webli 99.05 81.07 75.90
ViT-B-16-SigLIP-384 webli 99.40 84.42 80.98
ViT-B-16-SigLIP-512 webli 99.57 85.71 81.93
ViT-B-16-SigLIP-i18n-256 webli 99.48 89.41 86.14
ViT-B-16-plus-240 laion400m_e31 98.45 71.26 55.08
laion400m_e32 98.45 71.26 55.08
ViT-B-32 commonpool_m_basic_s128m_b4k 89.16 90.71 91.48
commonpool_m_clip_s128m_b4k 90.19 90.62 90.53
commonpool_m_image_s128m_b4k 89.07 91.31 92.69
commonpool_m_laion_s128m_b4k 87.26 90.10 90.02
commonpool_m_s128m_b4k 87.01 89.67 91.31
commonpool_m_text_s128m_b4k 88.55 90.96 91.22
commonpool_s_basic_s13m_b4k 72.03 74.27 74.96
commonpool_s_clip_s13m_b4k 76.51 76.42 76.33
commonpool_s_image_s13m_b4k 72.29 73.06 73.41
commonpool_s_laion_s13m_b4k 71.69 73.49 74.01
commonpool_s_s13m_b4k 73.32 74.01 73.84
commonpool_s_text_s13m_b4k 73.49 74.01 74.61
datacomp_m_s128m_b4k 89.16 90.62 91.05
datacomp_xl_s13b_b90k 98.11 78.66 65.40
laion2b_s34b_b79k 97.68 74.18 58.18
laion400m_e31 95.78 63.34 42.43
laion400m_e32 95.70 63.25 42.69
metaclip_400m 96.82 76.59 59.21
metaclip_fullcc 97.59 79.43 61.53
openai 96.82 60.50 35.89
ViT-B-32-256 datacomp_s34b_b86k 98.11 73.15 60.93
ViT-H-14 laion2b_s32b_b79k 98.88 63.08 49.66
metaclip_altogether 99.40 66.87 54.48
metaclip_fullcc 99.31 62.65 44.66
ViT-H-14-CLIPA datacomp1b 99.48 65.06 58.18
ViT-H-14-CLIPA-336 datacomp1b 99.57 61.62 56.54
laion2b 99.31 72.03 60.41
ViT-L-14 commonpool_xl_clip_s13b_b90k 99.05 57.75 50.09
commonpool_xl_laion_s13b_b90k 99.23 74.96 61.10
commonpool_xl_s13b_b90k 99.40 74.96 69.54
datacomp_xl_s13b_b90k 99.48 60.76 54.04
laion2b_s32b_b82k 98.80 62.91 50.77
laion400m_e31 98.97 67.38 53.01
laion400m_e32 98.97 66.95 53.36
metaclip_400m 98.45 58.00 41.65
metaclip_fullcc 99.14 63.43 46.39
openai 98.80 41.05 29.00
ViT-L-14-336 openai 99.23 34.68 27.45
ViT-L-14-CLIPA datacomp1b 99.40 72.81 64.63
ViT-L-14-CLIPA-336 datacomp1b 99.57 74.87 65.92
ViT-L-16-SigLIP-256 webli 99.40 86.75 81.24
ViT-L-16-SigLIP-384 webli 99.57 87.26 82.79
ViT-SO400M-14-SigLIP webli 99.57 84.51 83.30
ViT-SO400M-14-SigLIP-378 webli 99.66 85.80 84.94
ViT-SO400M-14-SigLIP-384 webli 99.66 85.63 84.94
ViT-SO400M-16-SigLIP-i18n-256 webli 99.40 88.81 84.17
ViT-bigG-14 laion2b_s39b_b160k 99.40 70.91 59.47
metaclip_fullcc 99.14 68.42 55.51
ViT-bigG-14-CLIPA datacomp1b 99.48 68.16 58.95
ViT-bigG-14-CLIPA-336 datacomp1b 99.57 64.03 57.14
ViT-g-14 laion2b_s12b_b42k 98.62 69.36 57.49
laion2b_s34b_b88k 98.71 62.05 48.02
convnext_base laion400m_s13b_b51k 98.11 58.69 41.22
convnext_base_w laion2b_s13b_b82k 98.71 66.52 55.08
laion2b_s13b_b82k_augreg 98.62 68.42 58.09
laion_aesthetic_s13b_b82k 98.11 67.56 54.48
convnext_base_w_320 laion_aesthetic_s13b_b82k 98.45 68.67 52.67
laion_aesthetic_s13b_b82k_augreg 98.36 71.94 61.27
convnext_large_d laion2b_s26b_b102k_augreg 99.23 67.30 56.02
convnext_large_d_320 laion2b_s29b_b131k_ft 99.05 67.04 58.09
laion2b_s29b_b131k_ft_soup 99.23 67.47 56.88
convnext_xxlarge laion2b_s34b_b82k_augreg 99.23 65.06 56.28
laion2b_s34b_b82k_augreg_rewind 99.05 66.27 55.34
laion2b_s34b_b82k_augreg_soup 99.23 65.15 55.85
nllb-clip-base v1 93.80 76.33 62.91
nllb-clip-base-siglip mrl 98.11 71.77 65.32
v1 98.36 77.54 76.68
nllb-clip-large v1 96.90 61.70 50.60
nllb-clip-large-siglip mrl 98.62 72.38 68.33
v1 99.05 74.01 76.16
roberta-ViT-B-32 laion2b_s12b_b32k 96.90 75.04 58.78
xlm-roberta-base-ViT-B-32 laion5b_s13b_b90k 97.85 76.25 60.76
xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 98.80 63.77 50.00
llava-llama3:8b - 97.83 40.17 43.88
llava:7b - 96.30 57.78 62.85
llava:13b - 97.68 62.42 56.66
llava:34b - 98.62 86.67 85.55
llama3.2-vision:90b - 99.05 70.75 72.33
gpt-4o-mini-2024-07-18 - 99.14 91.39 91.82
gpt-4o-2024-08-06 - 99.57 97.41 96.82
A.5 Architecture Comparison
In table 5 in the main text, we compare different architectures (CLIP-RN vs. CLIP-ViT, and CLIP-ViT vs. SigLIP-ViT) and state averaged accuracies. In table 10, we detail the specific models that we choose from the OpenCLIP suite for this comparison.
Table 10: Models chosen from the OpenCLIP suite for architecture comparison in table 5. The table groups models by architecture category and indicates the corresponding fixed training data. v1 refers to LAION-COCO NLLB [50].
Architecture Category Model Training Data
CLIP-ViT ViT-B-32 openai
ViT-B-16 openai
ViT-L-14 openai
ViT-L-14-336 openai
CLIP-RN RN50 openai
RN101 openai
RN50x4 openai
RN50x16 openai
RN50x64 openai
SigLIP-ViT nllb-clip-base-siglip v1
nllb-clip-large-siglip v1
CLIP-ViT nllb-clip-base v1
nllb-clip-large v1

### v3

@@
 Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems.
-We publicly release the datasets introduced in this paper under https://huggingface.co/datasets/BLISS-e-V/SCAM, along with the code for evaluations at https://github.com/Bliss-e-V/SCAM.
+We publicly release the datasets introduced in this paper along with the code for evaluations at www.bliss.berlin/research/scam.
 Figure 1: Overview of the SCAM datasets and benchmarking framework.
@@
 NoSCAM SCAM
-RN50 openai 97.50 36.32 ↓61.18
-ViT-B-32 laion2b 97.68 74.18 ↓23.50
-ViT-B-16 laion2b 98.54 68.93 ↓29.61
-ViT-B-16-SigLIP webli 99.05 81.50 ↓17.55
-ViT-L-14 commonpool_xl 99.40 74.96 ↓24.44
-openai 98.80 41.05 ↓57.75
-ViT-L-14-336 openai 99.23 34.68 ↓64.55
-ViT-g-14 laion2b 98.71 62.05 ↓36.66
-ViT-bigG-14 laion2b 99.40 70.91 ↓28.49
-llava-llama3-1.1:8b - 97.83 40.17 ↓57.66
-llava-1.6:7b - 96.30 57.78 ↓38.52
-llava-1.6:13b - 97.86 62.42 ↓35.44
-llava-1.6:34b - 98.62 86.67 ↓11.95
-llama3.2-vision:90b - 99.05 70.75 ↓28.30
-gpt-4o-mini-2024-07-18 - 99.14 91.39 ↓7.75
-gpt-4o-2024-08-06 - 99.57 97.41 ↓2.16
+RN50 openai 97.76 36.61 ↓61.15
+ViT-B-32 laion2b 98.45 74.68 ↓23.77
+ViT-B-16 laion2b 98.71 69.16 ↓29.55
+ViT-B-16-SigLIP webli 99.22 81.40 ↓17.82
+ViT-L-14 commonpool_xl 99.48 74.68 ↓24.80
+openai 99.14 40.14 ↓59.00
+ViT-L-14-336 openai 99.22 33.85 ↓66.37
+ViT-g-14 laion2b 99.05 61.93 ↓37.12
+ViT-bigG-14 laion2b 99.40 70.89 ↓28.51
+llava-llama3-1.1:8b - 98.09 39.50 ↓58.59
+llava-1.6:7b - 97.50 58.43 ↓39.07
+llava-1.6:13b - 98.88 58.00 ↓40.88
+llava-1.6:34b - 98.97 84.85 ↓14.11
+llama3.2-vision:90b - 98.88 71.01 ↓27.87
+gpt-4o-mini-2024-07-18 - 99.40 84.68 ↓14.72
+gpt-4o-2024-08-06 - 99.48 96.82 02.67
 2 Related Work
@@
 1A.
-To create the cleaned NoSCAM dataset, we remove all attack words by obscuring the post-it notes while preserving the original object images.
-Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font) and matching the colors of the post-its and ink from the original images.
+To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images.
+Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images.
 Data Annotation and Processing.
@@
 However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets.
-Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 35 percentage points.
+Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points.
 For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
@@
 Dataset Accuracy %
-commonpool_l_text 91.57
-commonpool_l_image 90.71
-commonpool_l_basic 89.76
-commonpool_l 88.47
-datacomp_l 79.60
-commonpool_l_laion 78.83
-commonpool_l_clip 70.91
-laion2b 68.93
-laion400m 68.80
-datacomp_xl 68.07
+commonpool_l_text 91.82
+commonpool_l_image 90.96
+commonpool_l_basic 90.09
+commonpool_l 88.72
+datacomp_l 80.10
+commonpool_l_laion 79.24
+commonpool_l_clip 71.15
+laion2b 69.16
+laion400m 69.04
+datacomp_xl 68.39
 Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to as v1).
@@
 Training data Architecture Accuracy %
-openai CLIP-ViT 44.06
-CLIP-RN 35.92
-v1 SigLIP-ViT 75.77
-CLIP-ViT 69.02
+openai CLIP-ViT 43.63
+CLIP-RN 35.76
+v1 SigLIP-ViT 75.80
+CLIP-ViT 69.04
 While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance.
@@
 Model Patch Size Accuarcy %
-ViT-B-32 32 74.18
-ViT-bigG-14 14 70.91
-ViT-B-16 16 68.93
-ViT-H-14 14 63.08
-ViT-L-14 14 62.91
-ViT-g-14 14 62.05
+ViT-B-32 32 74.68
+ViT-bigG-14 14 70.89
+ViT-B-16 16 69.16
+ViT-H-14 14 62.96
+ViT-L-14 14 62.53
+ViT-g-14 14 61.93
 Lastly, fig.
@@
 Figure 5: Susceptibility to typographic attack (evaluation on SCAM) is agnostic of VLMs size, measured in millions of parameters.
+Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean ± standard deviation across all 99 VLMs that we evaluate.
 Impact of the Size of the Attack.
@@
 The mean accuracy across all models follows this trend, which is consistent with the findings in [34].
-Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean ± standard deviation across all 99 VLMs that we evaluate.
 Prompt-Based Classification Results for LVLMs.
 Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility.
-Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 30-50 percentage points under typographic attacks.
+Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks.
 This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig.
@@
 This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions.
-In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 2 percentage points.
-The smaller gpt-4o-mini, however, experiences a higher drop of approximately 8 points, illustrating again that larger models tend to be more robust.
+In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 3 percentage points.
+The smaller gpt-4o-mini, however, experiences a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust.
 Despite that, model size alone does not determine robustness.
@@
 JH focused on related work research and supported overall writing.
+JL supported model evaluations, managed data and code hosting, and created the project webpage.
 LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback.
@@
 A.4 Model Accuracies
-Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM datasets.
+Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM, PAINT [38], and RTA-100 [36] datasets.
 Model Training data Accuracy (%)
-NoSCAM SCAM SynthSCAM
-RN101 openai 97.42 40.02 27.71
-RN50 openai 97.50 36.32 24.27
-RN50x16 openai 98.62 37.69 35.28
-RN50x4 openai 97.85 33.05 29.69
-RN50x64 openai 99.05 32.53 32.70
-ViT-B-16 commonpool_l_basic_s1b_b8k 98.54 89.76 82.87
-commonpool_l_clip_s1b_b8k 98.11 70.91 63.08
-commonpool_l_image_s1b_b8k 98.19 90.71 86.14
-commonpool_l_laion_s1b_b8k 97.42 78.83 67.30
-commonpool_l_s1b_b8k 97.50 88.47 81.33
-commonpool_l_text_s1b_b8k 98.02 91.57 85.97
-datacomp_l_s1b_b8k 98.02 79.60 69.62
-datacomp_xl_s13b_b90k 98.80 68.07 55.34
-laion2b_s34b_b88k 98.54 68.93 52.75
-laion400m_e31 98.02 68.93 52.93
-laion400m_e32 98.11 68.67 53.10
-metaclip_400m 97.85 58.78 37.95
-metaclip_fullcc 98.71 61.10 42.17
-openai 97.07 40.02 22.03
-ViT-B-16-SigLIP webli 99.05 81.50 76.25
-ViT-B-16-SigLIP-256 webli 99.05 81.07 75.90
-ViT-B-16-SigLIP-384 webli 99.40 84.42 80.98
-ViT-B-16-SigLIP-512 webli 99.57 85.71 81.93
-ViT-B-16-SigLIP-i18n-256 webli 99.48 89.41 86.14
-ViT-B-16-plus-240 laion400m_e31 98.45 71.26 55.08
-laion400m_e32 98.45 71.26 55.08
-ViT-B-32 commonpool_m_basic_s128m_b4k 89.16 90.71 91.48
-commonpool_m_clip_s128m_b4k 90.19 90.62 90.53
-commonpool_m_image_s128m_b4k 89.07 91.31 92.69
-commonpool_m_laion_s128m_b4k 87.26 90.10 90.02
-commonpool_m_s128m_b4k 87.01 89.67 91.31
-commonpool_m_text_s128m_b4k 88.55 90.96 91.22
-commonpool_s_basic_s13m_b4k 72.03 74.27 74.96
-commonpool_s_clip_s13m_b4k 76.51 76.42 76.33
-commonpool_s_image_s13m_b4k 72.29 73.06 73.41
-commonpool_s_laion_s13m_b4k 71.69 73.49 74.01
-commonpool_s_s13m_b4k 73.32 74.01 73.84
-commonpool_s_text_s13m_b4k 73.49 74.01 74.61
-datacomp_m_s128m_b4k 89.16 90.62 91.05
-datacomp_xl_s13b_b90k 98.11 78.66 65.40
-laion2b_s34b_b79k 97.68 74.18 58.18
-laion400m_e31 95.78 63.34 42.43
-laion400m_e32 95.70 63.25 42.69
-metaclip_400m 96.82 76.59 59.21
-metaclip_fullcc 97.59 79.43 61.53
-openai 96.82 60.50 35.89
-ViT-B-32-256 datacomp_s34b_b86k 98.11 73.15 60.93
-ViT-H-14 laion2b_s32b_b79k 98.88 63.08 49.66
-metaclip_altogether 99.40 66.87 54.48
-metaclip_fullcc 99.31 62.65 44.66
-ViT-H-14-CLIPA datacomp1b 99.48 65.06 58.18
-ViT-H-14-CLIPA-336 datacomp1b 99.57 61.62 56.54
-laion2b 99.31 72.03 60.41
-ViT-L-14 commonpool_xl_clip_s13b_b90k 99.05 57.75 50.09
-commonpool_xl_laion_s13b_b90k 99.23 74.96 61.10
-commonpool_xl_s13b_b90k 99.40 74.96 69.54
-datacomp_xl_s13b_b90k 99.48 60.76 54.04
-laion2b_s32b_b82k 98.80 62.91 50.77
-laion400m_e31 98.97 67.38 53.01
-laion400m_e32 98.97 66.95 53.36
-metaclip_400m 98.45 58.00 41.65
-metaclip_fullcc 99.14 63.43 46.39
-openai 98.80 41.05 29.00
-ViT-L-14-336 openai 99.23 34.68 27.45
-ViT-L-14-CLIPA datacomp1b 99.40 72.81 64.63
-ViT-L-14-CLIPA-336 datacomp1b 99.57 74.87 65.92
-ViT-L-16-SigLIP-256 webli 99.40 86.75 81.24
-ViT-L-16-SigLIP-384 webli 99.57 87.26 82.79
-ViT-SO400M-14-SigLIP webli 99.57 84.51 83.30
-ViT-SO400M-14-SigLIP-378 webli 99.66 85.80 84.94
-ViT-SO400M-14-SigLIP-384 webli 99.66 85.63 84.94
-ViT-SO400M-16-SigLIP-i18n-256 webli 99.40 88.81 84.17
-ViT-bigG-14 laion2b_s39b_b160k 99.40 70.91 59.47
-metaclip_fullcc 99.14 68.42 55.51
-ViT-bigG-14-CLIPA datacomp1b 99.48 68.16 58.95
-ViT-bigG-14-CLIPA-336 datacomp1b 99.57 64.03 57.14
-ViT-g-14 laion2b_s12b_b42k 98.62 69.36 57.49
-laion2b_s34b_b88k 98.71 62.05 48.02
-convnext_base laion400m_s13b_b51k 98.11 58.69 41.22
-convnext_base_w laion2b_s13b_b82k 98.71 66.52 55.08
-laion2b_s13b_b82k_augreg 98.62 68.42 58.09
-laion_aesthetic_s13b_b82k 98.11 67.56 54.48
-convnext_base_w_320 laion_aesthetic_s13b_b82k 98.45 68.67 52.67
-laion_aesthetic_s13b_b82k_augreg 98.36 71.94 61.27
-convnext_large_d laion2b_s26b_b102k_augreg 99.23 67.30 56.02
-convnext_large_d_320 laion2b_s29b_b131k_ft 99.05 67.04 58.09
-laion2b_s29b_b131k_ft_soup 99.23 67.47 56.88
-convnext_xxlarge laion2b_s34b_b82k_augreg 99.23 65.06 56.28
-laion2b_s34b_b82k_augreg_rewind 99.05 66.27 55.34
-laion2b_s34b_b82k_augreg_soup 99.23 65.15 55.85
-nllb-clip-base v1 93.80 76.33 62.91
-nllb-clip-base-siglip mrl 98.11 71.77 65.32
-v1 98.36 77.54 76.68
-nllb-clip-large v1 96.90 61.70 50.60
-nllb-clip-large-siglip mrl 98.62 72.38 68.33
-v1 99.05 74.01 76.16
-roberta-ViT-B-32 laion2b_s12b_b32k 96.90 75.04 58.78
-xlm-roberta-base-ViT-B-32 laion5b_s13b_b90k 97.85 76.25 60.76
-xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 98.80 63.77 50.00
-llava-llama3:8b - 97.83 40.17 43.88
-llava:7b - 96.30 57.78 62.85
-llava:13b - 97.68 62.42 56.66
-llava:34b - 98.62 86.67 85.55
-llama3.2-vision:90b - 99.05 70.75 72.33
-gpt-4o-mini-2024-07-18 - 99.14 91.39 91.82
-gpt-4o-2024-08-06 - 99.57 97.41 96.82
+NoSCAM SCAM SynthSCAM PAINT RTA-100
+RN101 openai 97.93 40.14 26.18 36.36 35.10
+RN50 openai 97.76 36.61 22.39 33.64 30.40
+RN50x16 openai 98.79 37.12 34.19 46.36 32.70
+RN50x4 openai 98.19 32.90 26.36 35.45 30.50
+RN50x64 openai 98.97 32.04 31.78 43.64 35.40
+ViT-B-16 commonpool_l_basic_s1b_b8k 98.62 90.09 83.03 82.73 88.10
+commonpool_l_clip_s1b_b8k 98.54 71.15 62.53 69.09 71.50
+commonpool_l_image_s1b_b8k 98.45 90.96 85.96 88.18 89.00
+commonpool_l_laion_s1b_b8k 97.59 79.24 65.72 77.27 72.90
+commonpool_l_s1b_b8k 98.02 88.72 81.65 86.36 87.20
+commonpool_l_text_s1b_b8k 98.36 91.82 85.10 85.45 90.00
+datacomp_l_s1b_b8k 98.19 80.10 69.51 71.82 76.60
+datacomp_xl_s13b_b90k 99.14 68.39 53.14 62.73 66.80
+laion2b_s34b_b88k 98.71 69.16 52.28 64.55 64.20
+laion400m_e31 98.62 69.16 50.90 66.36 62.60
+laion400m_e32 98.71 68.91 50.90 65.45 61.90
+metaclip_400m 98.62 58.66 35.40 50.00 57.10
+metaclip_fullcc 98.88 60.81 39.62 49.09 59.20
+openai 97.59 39.71 20.24 33.64 37.60
+ViT-B-16-SigLIP webli 99.22 81.40 75.45 82.73 78.10
+ViT-B-16-SigLIP-256 webli 99.31 80.96 75.97 83.64 80.00
+ViT-B-16-SigLIP-384 webli 99.48 84.24 80.45 84.55 82.20
+ViT-B-16-SigLIP-512 webli 99.66 85.70 81.14 83.64 82.90
+ViT-B-16-SigLIP-i18n-256 webli 99.40 89.66 85.70 89.09 85.70
+ViT-B-16-plus-240 laion400m_e31 98.97 71.75 54.44 63.64 62.70
+laion400m_e32 98.97 71.75 54.44 63.64 62.70
+ViT-B-32 commonpool_m_basic_s128m_b4k 89.15 90.78 91.47 92.73 92.70
+commonpool_m_clip_s128m_b4k 91.30 90.70 90.53 94.55 92.20
+commonpool_m_image_s128m_b4k 89.66 91.30 92.08 94.55 92.40
+commonpool_m_laion_s128m_b4k 89.23 90.35 90.18 94.55 91.20
+commonpool_m_s128m_b4k 88.11 89.84 90.70 95.45 90.60
+commonpool_m_text_s128m_b4k 89.32 91.04 90.96 94.55 93.40
+commonpool_s_basic_s13m_b4k 72.70 74.33 74.76 72.73 74.10
+commonpool_s_clip_s13m_b4k 76.31 76.57 76.66 76.36 77.10
+commonpool_s_image_s13m_b4k 72.95 73.21 73.47 78.18 72.50
+commonpool_s_laion_s13m_b4k 73.73 73.39 74.25 78.18 74.00
+commonpool_s_s13m_b4k 74.68 74.25 73.99 70.91 72.40
+commonpool_s_text_s13m_b4k 74.42 73.99 74.50 79.09 74.50
+datacomp_m_s128m_b4k 89.75 90.61 90.61 94.55 92.20
+datacomp_xl_s13b_b90k 98.88 79.33 64.17 69.09 76.30
+laion2b_s34b_b79k 98.45 74.68 56.16 68.18 66.00
+laion400m_e31 96.64 63.57 40.48 55.45 57.70
+laion400m_e32 96.64 63.48 40.57 54.55 57.40
+metaclip_400m 97.16 77.00 55.99 68.18 77.10
+metaclip_fullcc 98.45 79.93 58.91 67.27 75.30
+openai 96.47 60.81 33.59 52.73 51.40
+ViT-B-32-256 datacomp_s34b_b86k 98.54 73.30 60.72 70.91 70.00
+ViT-H-14 laion2b_s32b_b79k 99.22 62.96 47.63 51.82 55.60
+metaclip_altogether 99.48 66.58 52.89 61.82 58.30
+metaclip_fullcc 99.40 62.27 43.41 53.64 55.60
+ViT-H-14-CLIPA datacomp1b 99.66 64.69 57.02 49.09 57.30
+ViT-H-14-CLIPA-336 datacomp1b 99.66 61.15 55.64 46.36 59.10
+laion2b 99.22 71.92 60.47 60.00 61.60
+ViT-L-14 commonpool_xl_clip_s13b_b90k 99.48 57.62 48.84 44.55 51.80
+commonpool_xl_laion_s13b_b90k 99.40 74.94 59.09 62.73 67.40
+commonpool_xl_s13b_b90k 99.48 74.68 68.56 56.36 65.80
+datacomp_xl_s13b_b90k 99.48 60.38 52.97 50.00 56.00
+laion2b_s32b_b82k 99.05 62.53 48.23 59.09 56.40
+laion400m_e31 99.05 67.18 51.42 63.64 59.50
+laion400m_e32 99.05 66.75 52.02 63.64 59.90
+metaclip_400m 98.71 57.71 39.62 50.00 54.30
+metaclip_fullcc 99.14 63.14 45.48 47.27 52.60
+openai 99.14 40.14 26.61 40.00 36.90
+ViT-L-14-336 openai 99.22 33.85 26.44 33.64 39.90
+ViT-L-14-CLIPA datacomp1b 99.48 72.61 63.57 63.64 66.80
+ViT-L-14-CLIPA-336 datacomp1b 99.57 74.76 65.81 63.64 68.50
+ViT-L-16-SigLIP-256 webli 99.57 86.48 80.10 81.82 79.60
+ViT-L-16-SigLIP-384 webli 99.66 87.25 81.65 76.36 77.40
+ViT-SO400M-14-SigLIP webli 99.74 84.07 82.17 84.55 81.10
+ViT-SO400M-14-SigLIP-378 webli 99.66 85.62 85.01 83.64 81.30
+ViT-SO400M-14-SigLIP-384 webli 99.66 85.44 84.58 81.82 81.20
+ViT-SO400M-16-SigLIP-i18n-256 webli 99.74 88.89 83.98 85.45 86.30
+ViT-bigG-14 laion2b_s39b_b160k 99.40 70.89 58.91 56.36 58.50
+metaclip_fullcc 99.31 68.04 53.66 60.00 58.30
+ViT-bigG-14-CLIPA datacomp1b 99.66 68.04 59.35 56.36 63.20
+ViT-bigG-14-CLIPA-336 datacomp1b 99.74 64.00 56.68 58.18 59.40
+ViT-g-14 laion2b_s12b_b42k 99.14 69.25 55.38 59.09 54.70
+laion2b_s34b_b88k 99.05 61.93 47.03 54.55 50.10
+convnext_base laion400m_s13b_b51k 98.54 59.09 40.40 49.09 54.10
+convnext_base_w laion2b_s13b_b82k 98.79 66.84 52.28 60.91 59.40
+laion2b_s13b_b82k_augreg 98.88 68.73 56.33 60.00 57.40
+laion_aesthetic_s13b_b82k 98.45 68.13 51.59 59.09 59.70
+convnext_base_w_320 laion_aesthetic_s13b_b82k 98.62 68.91 51.25 55.45 61.40
+laion_aesthetic_s13b_b82k_augreg 98.62 72.27 59.00 59.09 63.90
+convnext_large_d laion2b_s26b_b102k_augreg 99.40 67.18 54.78 54.55 53.70
+convnext_large_d_320 laion2b_s29b_b131k_ft 99.31 67.10 57.11 55.45 56.90
+laion2b_s29b_b131k_ft_soup 99.57 67.53 56.16 53.64 57.00
+convnext_xxlarge laion2b_s34b_b82k_augreg 99.48 64.94 54.69 50.91 59.80
+laion2b_s34b_b82k_augreg_rewind 99.22 66.15 54.09 55.45 58.80
+laion2b_s34b_b82k_augreg_soup 99.48 64.94 54.44 55.45 59.00
+nllb-clip-base v1 94.66 76.66 61.67 71.82 74.90
+nllb-clip-base-siglip mrl 98.28 71.92 63.91 70.91 69.40
+v1 98.71 77.69 75.88 75.45 78.60
+nllb-clip-large v1 97.50 61.41 48.58 56.36 60.90
+nllb-clip-large-siglip mrl 98.88 72.27 66.75 66.36 71.20
+v1 99.05 73.90 75.37 74.55 75.90
+roberta-ViT-B-32 laion2b_s12b_b32k 97.67 75.45 57.45 62.73 68.50
+xlm-roberta-base-ViT-B-32 laion5b_s13b_b90k 98.62 76.83 59.00 69.09 71.20
+xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 99.22 63.74 47.89 54.55 56.70
+llava-llama3-1.1:8b - 98.09 39.50 44.04 48.62 41.83
+llava-1.6:7b - 97.50 57.25 61.45 53.64 51.20
+llava-1.6:13b - 98.88 58.00 55.42 54.55 56.00
+llava-1.6:34b - 98.97 84.85 85.11 89.09 81.50
+llama3.2-vision:90b - 98.88 71.01 74.65 77.06 68.41
+gpt-4o-mini-2024-07-18 - 99.40 84.68 87.09 90.91 85.15
+gpt-4o-2024-08-06 - 99.48 96.82 94.58 99.09 96.73
 A.5 Architecture Comparison

### v4

@@
 We publicly release the datasets introduced in this paper along with the code for evaluations at www.bliss.berlin/research/scam.
-Figure 1: Overview of the SCAM datasets and benchmarking framework.
-A) We create three counterfactual versions of images: SCAM, NoSCAM, and SynthSCAM.
-B) Evaluation of VLMs via cosine similarity between image and text.
-C) Prompt-based evaluation of LVLMs.
+Figure 1: SCAM enables benchmarking safety-critical robustness to typographic attacks in multimodal models.
+a) We construct three image variants.
+Real-world attacks in SCAM, a cleaned baseline NoSCAM, and digitally simulated attacks in SynthSCAM allow causal evaluation of typographic vulnerabilities.
+b) VLMs are evaluated zero-shot by computing cosine similarity between image embeddings and textual labels.
+Our results show that misleading text embedded in images significantly shifts predictions, indicating overreliance on textual cues.
+c) LVLMs are assessed using prompt-based classification.
+Despite advanced architectures, models remain susceptible to typographic attacks in realistic user-facing tasks.
+Notably, synthetic attacks are as effective as real ones, validating their use for scalable robustness evaluation.
 1 Introduction
-VLMs such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9].
+Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9].
 With the growing use of VLMs in real-world applications, such as medical applications [10, 11, 12] and autonomous systems [13], as well as in Large Vision-Language Models (LVLMs) [14, 15, 16], it is essential to ensure the safety and reliability of these models.
@@
 Data Annotation and Processing.
-All images were manually annotated with both object and corresponding attack labels as part of a voluntary community effort.
-The annotation process was conducted using Label Studio [39].
-To standardize the dataset, all images were zero-padded and downscaled to 512×512 pixels using cv2.INTER_AREA [40], an interpolation method that preserves details while minimizing distortion.
+[…]
 Dataset Statistics.
@@
 We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup.
-Both evaluations are conducted on the SCAM, NoSCAM and SynthSCAM datasets.
+Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT [38] and RTA-100 [36].
 The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
@@
 To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [41, 42, 1, 43, 44].
-We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 28, 29, 28] and Llama3.2-vision [45] models.
+We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 29, 28] and Llama3.2-vision [45] models.
 The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API.
@@
 (b) taxi
-Respond strictly with only (a) or (b), nothing else."
+Respond strictly with only (a) or (b), nothing else.
 […]
@@
 datacomp_xl 68.39
-Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to as v1).
-More details about the selection process can be in section A.5.
+Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1).
+More details about the selection process can be found in section A.5.
 Training data Architecture Accuracy %
@@
 Appendix A Appendix
-A.1 SCAM Examples
+A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
 In fig. 7, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
@@
 A.4 Model Accuracies
-Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM, PAINT [38], and RTA-100 [36] datasets.
+Table 9: Performance of VLMs and LVLMs available through OpenCLIP [44] resp.
+ollama [46] and OpenAI’s API on SCAM, PAINT [38], and RTA-100 [36] datasets.
+Results on PAINT and RTA-100 are broadly similar to those on SCAM, supporting the consistency of typographic attack effects across datasets.
 Model Training data Accuracy (%)
@@
 The table groups models by architecture category and indicates the corresponding fixed training data.
-v1 refers to LAION-COCO NLLB [50].
+In OpenCLIP, v1 refers to LAION-COCO NLLB [50].
 Architecture Category Model Training Data

### v6

@@
 Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images.
-However, existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities.
-In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1,162 images across hundreds of object categories and attack words.
-Through extensive benchmarking of Vision-Language Models (VLMs) on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks.
-Our findings reveal that typographic attacks persist in state-of-the-art Large Vision-Language Models (LVLMs) due to the choice of their vision encoder, though larger Large Language Models (LLMs) backbones help mitigate their vulnerability.
+Existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities.
+In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1162 images across hundreds of object categories and attack words.
+Through extensive benchmarking of Vision-Language Models on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks.
+Our findings indicate that typographic attacks remain effective against state-of-the-art Large Vision-Language Models, especially those employing vision encoders inherently vulnerable to such attacks.
+However, employing larger Large Language Model backbones reduces this vulnerability while simultaneously enhancing typographic understanding.
 Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research.
 Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems.
-We publicly release the datasets introduced in this paper along with the code for evaluations at www.bliss.berlin/research/scam.
+Finally, we publicly release the datasets introduced in this paper, along with the code for evaluations under www.bliss.berlin/research/scam.
+1 Introduction
+Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9].
+However, recent research has revealed that these models are vulnerable to typographic attacks [10, 11, 12], which exploit the reliance on the textual component of multimodal learning.
+Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions.
+Figure˜1 depicts how adding the word “taxi” to an image of a “clock” can cause a model to misclassify it as a “taxi”.
+These attacks expose a fundamental interaction in multimodal learning: the model gets tricked by focusing on text within an image, more than the visual content itself.
+Understanding and mitigating such vulnerabilities is essential for the safe deployment of multimodal foundation models.
 Figure 1: SCAM enables benchmarking safety-critical robustness to typographic attacks in multimodal models.
@@
 Notably, synthetic attacks are as effective as real ones, validating their use for scalable robustness evaluation.
-1 Introduction
-Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9].
 With the growing use of VLMs in real-world applications, such as medical applications [13, 14, 15] and autonomous systems [16], as well as in Large Vision-Language Models (LVLMs) [17, 18, 19], it is essential to ensure the safety and reliability of these models.
-However, recent research has revealed that these models are vulnerable to typographic attacks [17, 18, 19], which exploit the reliance on the textual component of multimodal learning.
-Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions.
-Figure 1 depicts how adding the word “taxi” to an image of a “clock” can cause a model to misclassify it as a “taxi”.
-These attacks expose a fundamental interaction in multimodal learning: the model gets tricked by focusing on text within an image, more than the visual content itself [17, 18, 19].
-Understanding and mitigating such vulnerabilities is essential for the safe deployment of multimodal foundation models.
-To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest real-world typographic attack dataset to date.
+To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date.
 Each image contains an object with a semantically unrelated handwritten attack word [^1] While we use the term “word”, our attacks also incorporate phrases like “no left turn”.
@@
 Furthermore, we provide two variations, a clean dataset where attack words are removed (NoSCAM), and a synthetic dataset where attack words are digitally reintroduced (SynthSCAM).
-We provide extensive evaluations over a range of VLMs and LVLMs and show that:
+In summary, we provide extensive evaluations over a range of VLMs and LVLMs and show that:
   1. 1.
-    Models of both classes experience a significant drop in accuracy when typographic attacks are introduced (SCAM).
+    Models of both classes experience a significant drop in accuracy when faced with real-world typographic attacks (SCAM).
   2. 2.
-    We find that synthetic attacks closely align with real-world attacks, providing empirical support for the established practice of evaluating on synthetic datasets.
+    Synthetic attacks (SynthSCAM) closely align with real-world attacks, providing empirical support for the established practice of evaluating on synthetic datasets.
   3. 3.
-    LVLMs, such as the LLaVA family, inherit typographic vulnerabilities from their vision encoders.
+    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses inherited from their vision encoders.
   4. 4.
-    Utilizing larger Large Language Model (LLM) backbones can alleviate an LVLM’s inherited weaknesses regarding typographic robustness.
-Table 1: Performance of VLMs and LVLMs available through OpenCLIP resp.
-ollama and OpenAI on the SCAM datasets.
+    Employing larger Large Language Model (LLM) backbones reduces susceptibility to typographic attacks in LVLMs while also improving their typographic understanding.
+Table 1: Performance of VLMs and LVLMs available through OpenCLIP [20] resp.
+ollama [21], and Anthropic’s resp.
+OpenAI’s API on the SCAM datasets.
 Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding.
 Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [22].
-This table presents a selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
+This table presents a only selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
 Model Training data Accuracy (%)
-NoSCAM SCAM
-RN50 openai 97.76 36.61 ↓61.15
-ViT-B-32 laion2b 98.45 74.68 ↓23.77
-ViT-B-16 laion2b 98.71 69.16 ↓29.55
-ViT-B-16-SigLIP webli 99.22 81.40 ↓17.82
-ViT-L-14 commonpool_xl 99.48 74.68 ↓24.80
-openai 99.14 40.14 ↓59.00
-ViT-L-14-336 openai 99.22 33.85 ↓66.37
-ViT-g-14 laion2b 99.05 61.93 ↓37.12
-ViT-bigG-14 laion2b 99.40 70.89 ↓28.51
-llava-llama3-1.1:8b - 98.09 39.50 ↓58.59
-llava-1.6:7b - 97.50 58.43 ↓39.07
-llava-1.6:13b - 98.88 58.00 ↓40.88
-llava-1.6:34b - 98.97 84.85 ↓14.11
-llama3.2-vision:90b - 98.88 71.01 ↓27.87
-gpt-4o-mini-2024-07-18 - 99.40 84.68 ↓14.72
-gpt-4o-2024-08-06 - 99.48 96.82 02.67
+NoSCAM SCAM SynthSCAM
+RN50 openai 97.76 36.61 $\downarrow$61.15 22.39 $\downarrow$75.37
+ViT-B-32 laion2b 98.45 74.68 $\downarrow$23.77 56.16 $\downarrow$42.29
+ViT-B-16 laion2b 98.71 69.16 $\downarrow$29.55 52.28 $\downarrow$46.43
+ViT-B-16-SigLIP webli 99.22 81.40 $\downarrow$17.82 75.45 $\downarrow$23.77
+ViT-L-14 commonpool_xl 99.48 74.68 $\downarrow$24.80 68.56 $\downarrow$30.92
+openai 99.14 40.14 $\downarrow$59.00 26.61 $\downarrow$72.53
+ViT-L-14-336 openai 99.22 33.85 $\downarrow$65.37 26.44 $\downarrow$72.78
+ViT-L-14-CLIPA-336 datacomp1b 99.57 74.76 $\downarrow$24.81 65.81 $\downarrow$33.76
+ViT-g-14 laion2b 99.05 61.93 $\downarrow$37.12 47.03 $\downarrow$52.02
+ViT-bigG-14 laion2b 99.40 70.89 $\downarrow$28.51 58.91 $\downarrow$40.49
+llava-llama3:8b - 98.09 39.50 $\downarrow$58.59 44.04 $\downarrow$54.05
+llava:7b-v1.6 - 97.50 58.43 $\downarrow$39.07 61.45 $\downarrow$36.06
+llava:13b-v1.6 - 98.88 58.00 $\downarrow$40.88 55.42 $\downarrow$43.46
+llava:34b-v1.6 - 98.97 84.85 $\downarrow$14.11 85.11 $\downarrow$13.86
+gemma3:4b - 97.24 58.05 $\downarrow$39.19 64.80 $\downarrow$32.44
+gemma3:12b - 99.14 52.02 $\downarrow$47.12 64.80 $\downarrow$34.34
+gemma3:27b - 97.42 81.67 $\downarrow$15.75 83.65 $\downarrow$13.77
+llama3.2-vision:90b - 98.88 71.01 $\downarrow$27.87 74.65 $\downarrow$24.22
+llama4:scout - 99.23 88.12 $\downarrow$11.10 87.18 $\downarrow$12.05
+gpt-4o-mini-2024-07-18 - 99.40 84.68 $\downarrow$14.72 87.09 $\downarrow$12.31
+claude-sonnet-4-20250514 - 99.31 91.13 $\downarrow$8.18 90.28 $\downarrow$9.03
+gpt-4o-2024-08-06 - 99.48 96.82 $\downarrow$2.67 94.58 $\downarrow$4.90
 2 Related Work
@@
 Following this paradigm, various Vision-Language Models (VLMs) [25, 26, 27] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7].
-Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [26, 27, 28, 29, 28].
+Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [28, 29, 30, 31, 32, 33, 34, 35].
 Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
@@
 Existing real-world datasets, such as PAINT [42] and Materzynska+[43], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models.
-The previously largest real-world typographic attack dataset, RTA-100 [36], consists of 1,000 images spanning 100 objects and attack words.
-However, existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see table 2 for a comparison).
+The previously largest real-world typographic attack dataset, RTA-100 [44], consists of 1000 images spanning 100 objects and attack words.
+However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison).
+Our dataset addresses this limitation by introducing substantially larger variability across object–word combinations, which mitigates dataset bias and enables a more reliable assessment of generalization in VLMs and LVLMs.
 Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains.
-This gap hinders not only the rigorous evaluation of VLM s and LVLM s but also the development of effective defense mechanisms.
-Figure 2: Examples of images from the SCAM dataset. Additional images for both NoSCAM and SynthSCAM are provided in fig. 7 in the appendix.
+SCAM is the first dataset to provide three explicitly aligned components: (i) SCAM, containing real-world attacked scenes, (ii) NoSCAM, the corresponding clean counterparts of the exact same scenes, and (iii) SynthSCAM, synthetically manipulated versions of the same object–word pairs.
+This tri-part design enables precise, controlled measurement of typographic attack effects and systematic comparisons across clean, real, and synthetic domains, establishing a stronger foundation for rigorous evaluation of VLMs and LVLMs as well as the development of effective defense mechanisms.
+Figure 2: Cropped examples from the SCAM dataset, illustrating the natural variability introduced by nine different contributors using distinct smartphones, handwriting styles, and recording environments.
+This diversity results in variations in lighting, camera quality, backgrounds, and writing appearance.
+Additional uncropped examples for both NoSCAM and SynthSCAM are provided in fig.˜7 in the appendix.
+3 SCAM Dataset
+We introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date, containing 1162 images (see LABEL:tab:dataset_comparison).
+The dataset and its two variations (NoSCAM and SynthSCAM), enabling empirical assessment of transferability between these settings, are publicly available on Hugging Face together with the fully compatible and publicly accessible SCAM benchmark repository for standardized evaluation.
+Data Collection.
+SCAM was collaboratively constructed by nine contributors using nice diffrent smartphones, ensuring natural variation in image quality.
+The data were recorded across diverse indoor and outdoor environments, including households, utility stores, and traffic infrastructure, under heterogeneous lighting conditions and viewing angles.
+This variability, illustrated in fig.˜2, reflects real-world conditions more faithfully than prior datasets.
+Each image in the SCAM dataset captures a physical object together with an unrelated attack word handwritten in English on a single post-it note and placed next to the object, as shown in fig.˜1a.
+Sticky notes were deliberately chosen as the attack medium because they are common in everyday contexts (e.g., labeling in offices, classrooms, and shops) and provide a realistic, reproducible surface for typographic attacks.
+To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images.
+This provides clean, paired counterparts for every attacked image.
+Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images.
+Choosing a uniform font style in this synthetic variant was a deliberate simplification.
+Prior synthetic attack datasets have already explored variations in font position, opacity, and color [40], whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
+Data Annotation and Processing.
+All images were manually annotated with object and attack word labels as well as post-it size (in %) as part of a voluntary community effort.
+The annotation process was conducted using Label Studio [45].
+To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA [46], an interpolation method that preserves details while minimizing distortion.
+Dataset Statistics.
+SCAM comprises 1162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations.
+The complete lists of objects and attack words are provided in section˜A.8.
+We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems.
+The distribution of these groups is depicted in fig.˜3.
 Figure 3: Distribution of attack words in SCAM into categories, highlighting both everyday terms and safety-critical vocabulary.
-3 SCAM Dataset
-We introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date, containing 1,162 images (see table 2).
-The dataset and its two variations (NoSCAM and SynthSCAM), allowing for an empirical assessment of transferability between these settings, are publicly available on Hugging Face, along with the fully compatible and publicly accessible SCAM benchmark repository for standardized evaluation.
-Data Collection.
-We collected 1,162 images from diverse environments, such as households, utility stores, and traffic infrastructure, ensuring a varied and representative dataset.
-An overview of example images from SCAM is presented in fig.
-2.
-Each image in the SCAM dataset captures an object together with an attack word – unrelated in meaning – handwritten on a single post-it note in English and placed alongside the object, as illustrated in fig.
-1a.
-To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images.
-Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images.
-Data Annotation and Processing.
-[…]
-Dataset Statistics.
-SCAM comprises 1,162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations.
-We provide lists of the 660 unique object labels and 206 attack words in section A.2.
-We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems.
-The distribution of these groups is depicted in fig.
-3.
-As shown in table 2, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words.
+Table 2: Comparison of real-world typographic attack datasets. SCAM is much more diverse in terms of objects and attack words.
+Dataset Images Objects Attacks
+PAINT [42] 110 89 87
+Materzynska+ [43] 180 20 8
+RTA-100 [44] 1000 100 100
+SCAM (ours) 1162 660 206
+As summarized in LABEL:tab:dataset_comparison, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words.
 Compared to RTA-100 [44], SCAM covers a much broader range of objects (660 vs.
@@
 100), offering a more comprehensive evaluation of model robustness to typographic attacks.
-This increased diversity makes SCAM a stronger benchmark for assessing the vulnerabilities of multimodal foundation models.
-Table 2: Comparison of real-world typographic attack datasets. SCAM is much more diverse in terms of objects and attack words.
-Dataset Images Objects Attack words
-PAINT [38] 110 89 87
-Materzynska+ [37] 180 20 8
-RTA-100 [36] 1,000 100 100
-SCAM (ours) 1,162 660 206
+The three explicitly aligned components – SCAM, NoSCAM, and SynthSCAM – allow precise, controlled measurement of typographic attack effects and direct comparisons between clean, real, and synthetic conditions, thereby closing critical gaps in prior datasets.
 4 Evaluation
@@
 To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [47, 48, 1, 49, 20].
-We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 29, 28] and Llama3.2-vision [45] models.
-The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API.
-An overview of selected models is provided in table 1, while LABEL:tab:full-overview presents the complete list of all 106 evaluated models.
+We also evaluate eleven LVLMs, including Claude Sonnet 4 [35], GPT-4o [28], LLaVA [29, 30, 31], Gemma3 [32], Llama3.2-vision [33] and Llama4 [34] models.
+The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform [21], whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp.
+Anthropic’s API.
+An overview of selected models is provided in table˜1, while LABEL:tab:full-overview presents the complete list of all 111 evaluated models.
 Evaluating VLMs.
-We evaluate the performance of VLMs in a zero-shot classification task on the SCAM datasets.
+We evaluate the performance of VLMs in a zero-shot classification task.
 For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word.
-These text embeddings are generated using the set of text templates proposed in [1], as detailed in section A.3.
+These text embeddings are generated using the set of text templates proposed in [1], as detailed in section˜A.2.
 The predicted label is then determined based on the highest cosine similarity score.
-An overview of this process is depicted in fig.
-1b.
+An overview of this process is depicted in fig.˜1b.
 Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly.
@@
 To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications.
-Our evaluation follows the setup illustrated in fig.
-1c, inspired by [47].
-The model is provided with an image from SCAM, NoSCAM, or SynthSCAM, along with the following prompt:
+Our evaluation follows the setup illustrated in fig.˜1c, inspired by [50].
+The model is provided with an image along with the following prompt:
 What entity is depicted in the image?
@@
 (b) taxi
-Respond strictly with only (a) or (b), nothing else.
+Respond strictly with only (a) or (b),
+nothing else.
+In this setup, the model correctly outputs “(a)” (the object label) when no attack word is present, as in the NoSCAM dataset.
+However, when the attack word is included, the model changes its response to “(b)”.
 […]
+All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and llama3.2-vision:90b and llama4:scout on an H100 80GB.
 5 Results
-We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table 1.
+We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table˜1.
 Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points.
-LVLMs exhibit similar behavior and we find that the choice of vision encoder plays a critical role in determining the extent of accuracy degradation.
-Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.
-4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table 3.
+LVLMs exhibit similar behavior, with the choice of vision encoder appearing to influence the extent of accuracy degradation.
+Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜4.
 These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
 LABEL:tab:full-overview in the appendix provides a full list of models and their accuracies.
-Table 3: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks.
-The tables present average confusion matrices.
-The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks.
-Values represent the mean ± standard deviation, rounded to integers.
-((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
-SynthSCAM SynthSCAM
-fails succeeds
-[…]
-[…]
-((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
-SynthSCAM SynthSCAM
-fails succeeds
-[…]
-[…]
-Zero-Shot Results for VLMs.
-As shown in fig.
-4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset.
-However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets.
-Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points.
-For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
-Impact of the Training Dataset and Model Architecture
-We investigate the vulnerability of VLMs by examining the impact of their training data, training paradigm, and model architecture, on their susceptibility to typographic attacks.
-Table 4 presents the impact of training data on the ViT-B models [48] trained by the OpenCLIP team.
-We focus solely on these models as they share the same training paradigm but are trained on different datasets.
-Overall, the ViT-B models trained on the LAION datasets [41] are highly susceptible to typographic attacks, while certain CommonPool variants [49] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively.
-Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI.
-Consequently, as indicated by the accuracies for that model in table 1, this filtering criterion appears to propagate susceptibility.
 Table 3: Training dataset choice significantly affects accuracy, as shown by the results on SCAM for the model ViT-B-16 trained by the OpenCLIP team on different datasets.
@@
 datacomp_xl 68.39
-Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1).
-More details about the selection process can be found in section A.5.
-Training data Architecture Accuracy %
+Table 4: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks.
+The tables present average confusion matrices.
+The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks.
+Values represent the mean $\pm$ standard deviation, rounded to integers.
+((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
+SynthSCAM SynthSCAM
+fails succeeds
+SCAM fails $648\pm 192$ $154\pm 68$
+SCAM succeeds $18\pm 11$ $341\pm 151$
+((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
+SynthSCAM SynthSCAM
+fails succeeds
+SCAM fails $741\pm 243$ $96\pm 47$
+SCAM succeeds $124\pm 71$ $201\pm 137$
+Zero-Shot Results for VLMs.
+As shown in fig.˜4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset.
+However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets.
+Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points.
+For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
+Impact of Training Dataset and Model Architecture.
+We investigate the vulnerability of VLMs by examining the impact of their training data, training paradigm, and model architecture, on their susceptibility to typographic attacks.
+Table 5: VLMs’ architecture impacts susceptibility to typographic attacks.
+SigLIP outperforms CLIP, ViT outperforms ResNets, and ViTs become more susceptible to typographic attacks as patch size decreases.
+The tables present (average) accuracy (Acc.) % on SCAM.
+((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin [51] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1).
+More details about the selection process can be found in section˜A.4.
+Training data Architecture Acc. %
 openai CLIP-ViT 43.63
@@
 CLIP-ViT 69.04
-While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance.
-We compare different architectures: CLIP-RN (ResNets [51]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [50] (see table 5).
-Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models.
-Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
-Our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table 6.
-Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
-Table 6: Models become more susceptible to typographic attacks as patch size decreases. Dataset fixed to laion2b, models evaluated on SCAM.
-Model Patch Size Accuarcy %
+((b)) Dataset fixed to laion2b.
+Model Patch Size Acc. %
 ViT-B-32 32 74.68
@@
 ViT-g-14 14 61.93
-Lastly, fig.
-6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks.
+Table˜3 presents the impact of training data on the ViT-B models [52] trained by the OpenCLIP team.
+We focus solely on these models as they share the same training paradigm but are trained on different datasets.
+Overall, the ViT-B models trained on the LAION datasets [47] are highly susceptible to typographic attacks, while certain CommonPool variants [53] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively.
+Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI.
+Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
+While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance.
+We compare different architectures: CLIP-RN (ResNets [54]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [51] (see table˜5(a)).
+Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models.
+Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
+Moreover, our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table˜5(b).
+Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
+Lastly, fig.˜6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks.
 However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
-Figure 5: Susceptibility to typographic attack (evaluation on SCAM) is agnostic of VLMs size, measured in millions of parameters.
-Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean ± standard deviation across all 99 VLMs that we evaluate.
 Impact of the Size of the Attack.
-We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.
-6).
+We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.˜6).
 The mean accuracy across all models follows this trend, which is consistent with the findings in [40].
+Figure 5: Susceptibility to typographic attack (accuracy on SCAM) is agnostic of VLM size, measured in millions of parameters.
+Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean $\pm$ standard deviation across all 99 VLMs that we evaluate.
 Prompt-Based Classification Results for LVLMs.
-Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility.
-Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks.
-This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig.
-6).
+Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table˜1) , their susceptibility appears to be influenced by the strength of the underlying vision encoder.
+Smaller LLaVA and Gemma3 models, such as gemma3:4b, llava:7b-v1.6, llava-llama3:8b, gemma3:12b, and llava:13b-v1.6, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks.
+We hypothesize that this vulnerability stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig.˜6).
 However, as model size increases, the robustness to typographic attacks improves.
-LLaVA models with larger LLM backbones, such as llava-1.6:34b, exhibit significantly better performance under attack.
+Models with larger LLM backbones, such as gemma:27b and llava:34b-v1.6, exhibit significantly better performance under attack.
 This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions.
-In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 3 percentage points.
-The smaller gpt-4o-mini, however, experiences a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust.
+In comparison, presumably much larger closed-source models such as claude-sonnet-4-20250514 and gpt-4o exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 8 resp.
+3 percentage points.
+The smaller gpt-4o-mini and llama4:scout, however, experience a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust.
 Despite that, model size alone does not determine robustness.
-In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava-1.6:34b, highlighting the critical role of the vision encoder in overall performance.
+In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava:34b-v1.6.
 These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
+Impact of VLM and LLM Size on LVLM Performance.
+Section˜A.6 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness.
+The experiment does not show a clear gain in robustness when substituting the vision encoder with one ranked safer by SCAM, while keeping the LLM, training data, and training regime unchanged.
+Instead, the overall performance decreases, highlighting the difficulty of choosing the optimal vision encoder and showcasing the need for further experiments investigating the robustness impact of different VLM choices.
+Apart from that, as discussed in section˜A.7, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.
 6 Discussion
@@
 Our analysis further reveals that training regime, training data, and model architecture significantly influence susceptibility of Vision-Language Models (VLMs) to typographic attacks, whereas model size alone does not correlate with robustness.
-In Large Vision-Language Models (LVLMs), vulnerability to typographic attacks carry over from the vision encoder, making these models inherently susceptible.
-However, we find that larger Large Language Model (LLM) backbones can mitigate this effect, leading to improved robustness and accuracy to typographic attacks.
+In Large Vision-Language Models (LVLMs), vulnerability to typographic attacks appears to carry over from the vision encoder, making these models susceptible to the same weaknesses.
+However, we find that larger Large Language Model (LLM) backbones can mitigate this effect, leading to improved robustness and accuracy to typographic attacks, while also improving their typographic understanding.
 Limitations.
-A key limitation of our work is the use of “naive prompting” [52, 53] for LVLMs.
-Utilizing multi-prompt evaluation, such as those proposed in [54, 47], could provide a better assessment of their robustness to typographic attacks.
-Another limitation is the focus on English-language attacks and it should be assessed whether language-specific factors influence model susceptibility.
+Despite incorporating a different, “safer” prompt in section˜A.5, a central limitation of our evaluation remains the reliance on “naive prompting” [55, 56] for assessing LVLMs.
+Adopting more advanced multi-prompt evaluation strategies may yield a more reliable measure of model robustness against typographic attacks.
+A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility.
 Future Work.
 As multimodal models are increasingly deployed to safety-critical applications [15, 14, 16, 16], ensuring robustness against typographic attacks is essential to prevent serious consequences.
-Approaches such as alignment training [55] or defense techniques [36] offer promising directions for enhancing robustness.
-Additionally, our findings suggest that an LVLM ’s vulnerability to typographic attacks is closely tied to its vision encoder.
-To investigate this further, we plan to retrain LLaVA models with different vision backbones to systematically evaluate their impact on robustness.
-This will help determine whether stronger vision encoders mitigate susceptibility or if additional fine-tuning strategies are required.
+Approaches such as alignment training [57] or (mechanistic) defense techniques [44, 58] offer promising directions for enhancing robustness.
+Our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder.
+However, our initial experiments (section˜A.6) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively.
+Future work will extend this by systematically retraining LLaVA models with a range of vision backbones to evaluate their role in attack susceptibility and to assess whether encoder strength alone suffices or if additional adaptation steps are necessary.
 7 Conclusion
@@
 Our evaluation demonstrated that current state-of-the-art multimodal models are significantly affected by typographic attacks, highlighting a critical weakness in their robustness.
-We hope that our broad analysis of Vision-Language Models (VLMs) will help researchers make more informed choices when selecting typographically robust models for their downstream applications.
+We expect that our broad analysis of Vision-Language Models (VLMs) and Large Vision-Language Models (LVLMs) will help researchers make more informed choices when selecting typographically robust models for their safety-critical downstream applications.
 Acknowledgments and Disclosure of Funding
@@
 We thank the BLISS [^2] https://bliss.berlin/ community – its very existence, collaborative energy, and shared commitment to open research made this work possible.
-Special thanks to Zeynep Altayli, Jonas Loos, Philippa Ringe, and Joseph Tschörner for their valuable efforts in image collection and labeling.
-We especially thank Jonas Loos for his significant work on the project page.
+Special thanks to Zeynep Altayli, Philippa Ringe, and Joseph Tschörner for their valuable efforts in image collection and labeling.
 Author Contributions
@@
 LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback.
+EH provided supervision and contributed through discussions.
 LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work.
-All authors contributed to image collection and labeling.
+All authors except EH contributed to image collection and labeling.
 Appendix A Appendix
 A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
-In fig. 7, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
+In fig.˜7, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
 Figure 7: Example images of SCAM, NoSCAM, and SynthSCAM.
-A.2 SCAM Labels
-The SCAM dataset includes diverse object labels:
-30kmh traffic zone sign, adapter, advent calendar, alarm clock, analog timer, apple, apple seed cutter, apples, armchair, ashtray, aubergine, avocado, axe, back, backpack, bacon bits, badminton ball, bag, baking pan, ball, balloon, banana, bananas, band-aid, barricade, baseball cap, basket, basket ball, bathroom sign, bathtub, batteries, bed, bell, bell pepper, bench, bench vise, bicycle, bicycle bell, bicycle light, bicycle tire, bike, bike light, bike lock, bin, bird feeder, blackboard, blanket, blender, blower, board, boat figurine, book, boot, bottle, bottle cap, bottle crate, bottle opener, bottles, boulder, bowl, box, boxes, brass tubes, bread slicer, bricks, broom, brooms, brush, brushes, bucket, butterfly, buttons, cabinet, cable, cable drum, cable ties, cage, cake, calculator, calendar, camera, candle, candles, candy cane, canister, cans, cap, capacitor, car, carafe, cardboard box, carrot, cash register, cashew nuts, cat figurine, cat tower, cauldron, celery, chain, chain saw, chair, chalk bag, champagne, charger, cheese, chess board, chess piece, chess set, chessboard, chewing gum, chewing gums, chili pepper, chips, chocolate, chopsticks, cigarettes, cleaner, clipboard, clock, closet, clothes iron, clothespin, coat, coat hanger, coat hook, coffee beans, coffee machine, coffee maker, coin, coin counter, combination lock, compass, condom, condoms, confetti, contact grill, controller, cooking scraper, cordless drill, cork, corn, counter, cow meat, crepe pan, crochet hook, cross wrench, cucumber, cup, deodorant, detergent, detergent powder, diapers, dice, diffuser, dish scrubber, dish soap, dishwasher, dishwasher tablets, disinfectant dispenser, dispenser, diving goggles, dog prohibition sign, doll, door, doorhandle, doormat, drill, drills, drink, drone, drum, duct tape, dustpan, dvd, e-scooter, ear phones, ear plugs, ear protection, easel, edamame, edding, eggs, electric kettle, electric screwdriver, electric shaver, electric toothbrush, electrical panel, elephant, emergency exit sign, end of parking zone sign, ernie, exit sign, faucet, feeder, fence, fennel, fever thermometer, figurine, fineliner pen, fire alarm button, fire detector, fire extinguisher, firework, first aid kit, fish, flip flops, flipper, floss, flower, flower pot, flowering pot, flowers, folder, folding rule, fork, fortune cookie, freeze pack, french press, fridge, fries, frisbee, frothing pitcher, fryer, funnel, garden gnome, garlic, gas bottle, gas cooker, gem lettuce, ginger, glas, glass, glass bottle, glasses, globe, glove, glue stick, gnocchi, goal, goggles, goose, grapes, grater, grave candles, green beans, grill, guitar, guitar foot stand, hair clip, hair clipper, hair dryer, hair iron, hairbrush, hairgrip, hairspray, ham, hammer, hand blender, hand brush and dustpan, handbrush, handkerchiefs, handwash, hanger, hard disk, hare, hat, hdmi splitter, headband, headphones, heater, helmet, high visibility jacket, hook, horn, hot dog maker, hot glue gun, hot water bottle, hotplate, ice cube tray, ice pack, ice scraper, icecream, intercom, interdental brush, iron, ironing board, jacket, jam, jar, jersey, juggling ball, juicer, kettle, key, key chain, keyboard, keys, kiwi, knife, labeling machine, ladder, lamp, landing net, lantern, laptop, lawn chair, letter box, lid, light bulbs, light switch, lightbulb, lighter, lime, lipstick, lollies, lollipop, loofah, macbook, magazine, makeup powder, mango, marker, masher, mask, matryoshka doll, meatball, melon, memory card, menstrual pad, menstrual pads, mercedes-benz, metronome, microcontroller, microphone, microwave, milk pitcher, mirror, moka pot, money, monitor, mosquito repellent, mosquito spray, motion detector, mouse, mug, mulled wine, multiple socket, mushroom, mushrooms, nail clipper, nail polish, napkin, necklace, no entry sign, no parking sign, no pedestrian crossing sign, no smoking sign, no-entry sign, noodle cooker, notebook, nuts, oat bar, oats, octopus, office chair, olives, onion, onions, orange, organizer, oven, oven glove, paint, paint brush, paint roller, painting, pair, palette knife, pallet, pan, panini press, pants, paper rolls, paper streamers, paper towel, paper towels, paper wrap, paprika, parking sign, parking ticket machine, parsley, party hat, pasta, pea, pen, pencil, pepper, pepper caster, perfume, person, phone, pick, pickaxe, piggy bank, pillow, pills, pine cone, pineapple plant, pineapples, ping pong ball, pipe connection, pipes, pizza, pizza cutter, pizza oven, pizza peel, plane ticket, plant, plastic cups, plate, playing card, plectrum, plug board, plug socket, plunger, pocket balance, pocket knife, pokemon ball, pole, portable electric drill, post, post-its, pot, potato, potato chips, potato masher, potatoes, pouch, pouffe, power drill, powerbank, printer, printer cartridge, projector, puffed rice, pump, pumpkin, puzzle, puzzle piece, rack, racket, radiator, radio, railing, rake, razor, record player, remote, remote control, resistor, ribbon, rice cooker, roll board, roll-on, roller, roller board, rolling pin, rope, rose, rosemary, roses, router, rubber bands, rubber duck, rubik’s cube, running shoes, safe, safety pin, safety pins, safety sign, sage leaves, salad, salami, salmon, salt, sandal, sandpaper, sandwich maker, santa figurine, sausage, sausages, saw, scissor, scissors, scooter, screen, screw, screwdriver, screws, seafood salad, serving tray, serving trolley, sewing machine, shaker, shampoo, shashlik skewers, shaver, shelf, shell, shirt, shoe, shoes, shopping cart, shovel, shovels, shower, shower gel, shower head, shrimp, shuttlecock, sieve, sign, silverware box, sink, skillet, skimmer, skimmers, sledgehammer, sleeping mat, slider, slipper, slippers, sneaker, soap, soap dispenser, sock, socket, socks, spade, speaker, spice mill, spinach, spiral notebooks, sponge, spoon, spout, spray bottle, spray can, spring, squeegee, stamp, stapler, steam flat iron, step bin, stereo, stone plates, stool, stop sign, storage box, straw, straws, string, stuffed animal, suitcase, sunglasses, sunscreen, sweeper, t-shirt, table, takeaway bag, tampon, tap, tape, tape dispenser, tea, tea pot, teabag, tealight, tee box, telephone, tennis ball, terminal, thermal camping mat, thermometer, thread, ticket machine, tin box, tire, tissues, toaster, toilet, toilet bowl, toilet brush, toilet lid, toilet paper, tomato, tomatoes, tongue brush, toothbrush, toothpaste, tortellini, towel, toy car, toy elephant, toy horse, traffic light, traffic signs, trailer, trash bag, trash bin, trash can, trashcan, tray, tree, try square, turmeric, turtle, tv, ufonauts, umbrella, underpants, usb cable, usb-c cable, vacuum, vacuum cleaner, vase, vegetable peeler, video surveillance sign, vinyl, waffles, wall, wall charger, wallet, washbasin, washing machine, watch, water wings, watering can, watermelon, webcam, weight, wet floor sign, wheel, whisk, white cabbage, wildlife camera, window, window cleaner, wine, wood, wooden disc, wooden grill tong, wooden panels, wooden pestle, wooden skewer, wool, wrapping paper, wrench, yarn, and zucchini.
-Table 7 summarizes the frequency of object labels.
-The top 15 labels (e.g., “lamp”, “plant”, “bottle”) are among the most frequent, while the bottom 5 labels (e.g., “chain”, “roller board”, “mask”) appear rarely.
-Notably, 437 object labels appear exactly once, illustrating a pronounced long-tail distribution that underscores the diversity of the dataset.
-Table 7: Top 15 and bottom 5 object labels by appearance frequency.
-Object label Count
-lamp 13
-plant 13
-bottle 13
-chair 11
-glass 10
-candle 10
-cable 8
-grill 8
-printer 8
-cup 8
-speaker 8
-pillow 8
-brush 7
-canister 7
-pen 7
-chain 1
-roller board 1
-mask 1
-goggles 1
-webcam 1
-SCAM includes a variety of attack words:
-apple, arrow, bag, ball, barrier, basketball, beach, beatles, bed, beer, bell, bench, bicycle, bike, bird, blanket, boat, book, boot, bottle, boulder, bowl, box, boy, bracelet, bridge, bridge out, broom, bus, butter, cable, cake, camera, canal, candle, cane, cannon, cap, car, carpet, carrot, cart, cat, cellphone, chair, cheese, chest, chicken, child, clock, cloud, coat, coin, cone, couch, cow, crab, crystal, cup, curb, curtaxin, curve, dead end, deer, desert, desk, detour, dish, dog, door, duck, eagle, egg, elevator, exit, fan, feather, fence, finger, fire, fish, flag, floor, flower, fog, forest, fork, frog, fruit, garbage bin, gears, ginger, gloves, grandpa, grape, grass, gravel, green, grenade, gun, hair, hammer, hat, helmet, highway, hill, hole, honey, horn, horse, house, ice, iron, jacket, jar, jeep, key, knife, ladder, lake, lamp, lane, leaf, lemon, lentils, lift, lightbulb, lion, lipstick, lock, macbook, mad, median, merge, mirror, mittens, mountaxin, no left turn, no parking, ocean, olive oil, oregano, pedestrian, pen, person, phone, piano, picasso, pig, police, pumpkin, raccoon, red light, risotto rice, river, road, road closed, rock, saddle, salt, sandwich, santa claus, scarf, scooter, screw, screwdriver, sea, seed, shadow, ship, sidewalk, signal, slippery, sloth, slow, smartphone, sofa, speed camera, speed limit, spok, spoon, staxirs, stop, stop sign, store, street, street sign, suv, table, taxi, toll, traffic light, traffic sign, train, transistor, trash, tree, truck, unicorn, van, vase, wall, wet, woman, yield, and zebra.
-Table 8 summarizes the frequency of attack words.
-The top 15 attack words (e.g., “knife”, “truck”, “slow”) occur frequently, whereas many attack words are rare.
-In fact, 56 attack words appear exactly once, again highlighting overall diversity of the dataset.
-Table 8: Top 15 and bottom 5 attack words by appearance frequency.
-Attack Word Count
-knife 24
-truck 23
-slow 20
-lion 20
-zebra 19
-helmet 18
-bike 17
-honey 15
-pig 15
-camera 15
-lemon 15
-pen 15
-tree 14
-house 14
-fish 13
-beatles 1
-green 1
-raccoon 1
-beach 1
-unicorn 1
-A.3 Zero-Shot Evaluation of LVMs: Text Templates
+A.2 Zero-Shot Evaluation of LVMs: Text Templates
 To compute the cosine similarity between an image and text embeddings in our zero-shot evaluation of Vision-Language Models (VLMs), we generate text embeddings for both the object label and the attack word using multiple textual variations.
@@
 The templates are as follows: ( ’a bad photo of a {}.’, ’a photo of many {}.’, ’a sculpture of a {}.’, ’a photo of the hard to see {}.’, ’a low resolution photo of the {}.’, ’a rendering of a {}.’, ’graffiti of a {}.’, ’a bad photo of the {}.’, ’a cropped photo of the {}.’, ’a tattoo of a {}.’, ’the embroidered {}.’, ’a photo of a hard to see {}.’, ’a bright photo of a {}.’, ’a photo of a clean {}.’, ’a photo of a dirty {}.’, ’a dark photo of the {}.’, ’a drawing of a {}.’, ’a photo of my {}.’, ’the plastic {}.’, ’a photo of the cool {}.’, ’a close-up photo of a {}.’, ’a black and white photo of the {}.’, ’a painting of the {}.’, ’a painting of a {}.’, ’a pixelated photo of the {}.’, ’a sculpture of the {}.’, ’a bright photo of the {}.’, ’a cropped photo of a {}.’, ’a plastic {}.’, ’a photo of the dirty {}.’, ’a jpeg corrupted photo of a {}.’, ’a blurry photo of the {}.’ ).
-A.4 Model Accuracies
-Table 9: Performance of VLMs and LVLMs available through OpenCLIP [44] resp.
-ollama [46] and OpenAI’s API on SCAM, PAINT [38], and RTA-100 [36] datasets.
-Results on PAINT and RTA-100 are broadly similar to those on SCAM, supporting the consistency of typographic attack effects across datasets.
+A.3 Model Accuracies on Typographic Attack Datasets
+Table 6: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on SCAM, PAINT [42], and RTA-100 [44] datasets.
 Model Training data Accuracy (%)
@@
 xlm-roberta-large-ViT-H-14 frozen_laion5b_s13b_b90k 99.22 63.74 47.89 54.55 56.70
-llava-llama3-1.1:8b - 98.09 39.50 44.04 48.62 41.83
-llava-1.6:7b - 97.50 57.25 61.45 53.64 51.20
-llava-1.6:13b - 98.88 58.00 55.42 54.55 56.00
-llava-1.6:34b - 98.97 84.85 85.11 89.09 81.50
+llava-llama3:8b - 98.09 39.50 44.04 48.62 41.83
+llava:7b-v1.6 - 97.50 58.43 61.45 53.64 51.20
+llava:13b-v1.6 - 98.88 58.00 55.42 54.55 56.00
+llava:34b-v1.6 - 98.97 84.85 85.11 89.09 81.50
+gemma3:4b - 97.24 58.05 64.80 50.00 60.66
+gemma3:12b - 99.14 52.02 64.80 44.55 50.50
+gemma3:27b - 97.42 81.67 83.65 80.91 78.10
 llama3.2-vision:90b - 98.88 71.01 74.65 77.06 68.41
+llama4:scout - 99.23 88.12 87.18 89.09 86.20
 gpt-4o-mini-2024-07-18 - 99.40 84.68 87.09 90.91 85.15
+claude-sonnet-4-20250514 - 99.31 91.13 90.28 95.45 87.30
 gpt-4o-2024-08-06 - 99.48 96.82 94.58 99.09 96.73
-A.5 Architecture Comparison
-In table 5 in the main text, we compare different architectures (CLIP-RN vs.
+A.4 Architecture Comparison
+In table˜5(a) in the main text, we compare different architectures (CLIP-RN vs.
 CLIP-ViT, and CLIP-ViT vs.
-SigLIP-ViT) and state averaged accuracies.
-In table 10, we detail the specific models that we choose from the OpenCLIP suite for this comparison.
-Table 10: Models chosen from the OpenCLIP suite for architecture comparison in table 5.
+SigLIP-ViT) and present averaged accuracies.
+Table˜7 details the specific models that we choose from the OpenCLIP suite for this comparison.
+Table 7: Models chosen from the OpenCLIP suite for architecture comparison in table˜5(a).
 The table groups models by architecture category and indicates the corresponding fixed training data.
@@
 nllb-clip-large v1
+A.5 Prompt Variation
+Our evaluation of LVLMs in section˜4 focuses on a single prompt.
+Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation [59, 50], we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt.
+Contrary to the expectation formulated in these works, this modification does not reduce vulnerability.
+As illustrated in fig.˜8, the susceptibility to typographic attacks remains largely unchanged.
+Figure 8: The modified “safer” prompt does not alleviate susceptibility to typographic attacks.
+A.6 Replacing VLM in LLaVA
+Previous work [40, 60] suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact.
+Further, [61] demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in [60].
+Our evaluations on SCAM support the conclusion that some VLMs exhibit higher typographic robustness than others.
+In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA [14, 30, 31] – show heightened vulnerability.
+To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder [62, 63], which was trained on datacomp1b and differs primarily in training data and regime, not architecture.
+Compared to the OpenAI encoder, the CLIPA model exhibits a substantially lower robustness drop: only 25% from NoSCAM to SCAM, versus 65% (see table˜1).
+It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP [48] [^3] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
+[…]
+Table˜8 shows that our CLIPA-based LLaVA variant does not outperform the original model on SCAM, despite using a VLM that is more robust in isolation.
+Performance on other benchmarks evaluated via VLMEvalKit [65] also drops.
+These findings suggest that the benefits of a more robust vision encoder do not transfer under naive LLaVA integration.
+Note the experimental constraints: only one small Large Language Model (LLM) (Vicuna v1.5 7B), only two different vision backbones, and no hyperparameter tuning are employed.
+Table 8: Performance comparison of the original LLaVA-v1.5-7B and our CLIPA-based LLaVA variant across diverse benchmarks.
+Despite the CLIPA vision encoder showing stronger robustness in isolation, this does not translate to improved performance on SCAM.
+Evaluations for the upper section are conducted using VLMEvalKit [65].
+Other results follow the protocol outlined in the main part of the paper.
+Benchmark LLaVA-v1.5-7B (original) LLaVA-v1.5-7B-CLIPA (ours)
+MMBench v1.1 [66] 64.32 60.22
+MME Perception [67] 1444.81 1378.96
+SEED-Bench-1 (img) [68] 66.81 64.02
+TextVQA [69] 46.93 45.57
+OCRBench [70] 31.90 31.60
+OCR-VQA [71] 60.64 57.39
+NoSCAM 97.16 89.66
+SCAM 54.69 45.48
+SynthSCAM 51.55 46.08
+PAINT [42] 52.73 53.64
+RTA100 [44] 48.30 47.30
+A.7 Typographic Robustness vs. Typographic Understanding
+The main paper shows that increasing the size of LLM backbones reduces susceptibility to typographic attacks (see results for LLaVA and Gemma3 in table˜1).
+Simultaneously, benchmarks such as TextVQA and OCRBench, reported on the VLMLeaderboards [^5] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, indicate that typographic understanding – essentially, reading – also improves with model scale (see table˜9).
+Together, these findings suggest that typographic robustness emerges as model size increases, alongside improved textual perception.
+However, [36] reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
+Table 9: Performance of selected LLaVA-Next [31] models on three OCR-related benchmarks (TextVQA [69], OCRBench [70] (from VLMEvalKit [65]), and OCR-VQA [71]) shows that increasing LLM backbone size leads to improved typographic understanding.
+All models use the ViT-L-14-336 vision encoder.
+Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit [65].
+Model Params (B) TextVQA (%) OCRBench OCR-VQA (%)
+LLaVA-Next-Yi-34B 34.8 69.3 574 66.3
+LLaVA-Next-Vicuna-13B 13.4 66.9 537 65.3
+LLaVA-Next-Llama3 8.0 65.3 531 60.7
+LLaVA-Next-Mistral-7B 7.6 65.2 531 61.2
+LLaVA-Next-Vicuna-7B 7.1 63.8 532 63.8
+A.8 SCAM Labels
+The SCAM dataset includes diverse object labels:
+30kmh traffic zone sign, adapter, advent calendar, alarm clock, analog timer, apple, apple seed cutter, apples, armchair, ashtray, aubergine, avocado, axe, back, backpack, bacon bits, badminton ball, bag, baking pan, ball, balloon, banana, bananas, band-aid, barricade, baseball cap, basket, basket ball, bathroom sign, bathtub, batteries, bed, bell, bell pepper, bench, bench vise, bicycle, bicycle bell, bicycle light, bicycle tire, bike, bike light, bike lock, bin, bird feeder, blackboard, blanket, blender, blower, board, boat figurine, book, boot, bottle, bottle cap, bottle crate, bottle opener, bottles, boulder, bowl, box, boxes, brass tubes, bread slicer, bricks, broom, brooms, brush, brushes, bucket, butterfly, buttons, cabinet, cable, cable drum, cable ties, cage, cake, calculator, calendar, camera, candle, candles, candy cane, canister, cans, cap, capacitor, car, carafe, cardboard box, carrot, cash register, cashew nuts, cat figurine, cat tower, cauldron, celery, chain, chain saw, chair, chalk bag, champagne, charger, cheese, chess board, chess piece, chess set, chessboard, chewing gum, chewing gums, chili pepper, chips, chocolate, chopsticks, cigarettes, cleaner, clipboard, clock, closet, clothes iron, clothespin, coat, coat hanger, coat hook, coffee beans, coffee machine, coffee maker, coin, coin counter, combination lock, compass, condom, condoms, confetti, contact grill, controller, cooking scraper, cordless drill, cork, corn, counter, cow meat, crepe pan, crochet hook, cross wrench, cucumber, cup, deodorant, detergent, detergent powder, diapers, dice, diffuser, dish scrubber, dish soap, dishwasher, dishwasher tablets, disinfectant dispenser, dispenser, diving goggles, dog prohibition sign, doll, door, doorhandle, doormat, drill, drills, drink, drone, drum, duct tape, dustpan, dvd, e-scooter, ear phones, ear plugs, ear protection, easel, edamame, edding, eggs, electric kettle, electric screwdriver, electric shaver, electric toothbrush, electrical panel, elephant, emergency exit sign, end of parking zone sign, ernie, exit sign, faucet, feeder, fence, fennel, fever thermometer, figurine, fineliner pen, fire alarm button, fire detector, fire extinguisher, firework, first aid kit, fish, flip flops, flipper, floss, flower, flower pot, flowering pot, flowers, folder, folding rule, fork, fortune cookie, freeze pack, french press, fridge, fries, frisbee, frothing pitcher, fryer, funnel, garden gnome, garlic, gas bottle, gas cooker, gem lettuce, ginger, glas, glass, glass bottle, glasses, globe, glove, glue stick, gnocchi, goal, goggles, goose, grapes, grater, grave candles, green beans, grill, guitar, guitar foot stand, hair clip, hair clipper, hair dryer, hair iron, hairbrush, hairgrip, hairspray, ham, hammer, hand blender, hand brush and dustpan, handbrush, handkerchiefs, handwash, hanger, hard disk, hare, hat, hdmi splitter, headband, headphones, heater, helmet, high visibility jacket, hook, horn, hot dog maker, hot glue gun, hot water bottle, hotplate, ice cube tray, ice pack, ice scraper, icecream, intercom, interdental brush, iron, ironing board, jacket, jam, jar, jersey, juggling ball, juicer, kettle, key, key chain, keyboard, keys, kiwi, knife, labeling machine, ladder, lamp, landing net, lantern, laptop, lawn chair, letter box, lid, light bulbs, light switch, lightbulb, lighter, lime, lipstick, lollies, lollipop, loofah, macbook, magazine, makeup powder, mango, marker, masher, mask, matryoshka doll, meatball, melon, memory card, menstrual pad, menstrual pads, mercedes-benz, metronome, microcontroller, microphone, microwave, milk pitcher, mirror, moka pot, money, monitor, mosquito repellent, mosquito spray, motion detector, mouse, mug, mulled wine, multiple socket, mushroom, mushrooms, nail clipper, nail polish, napkin, necklace, no entry sign, no parking sign, no pedestrian crossing sign, no smoking sign, no-entry sign, noodle cooker, notebook, nuts, oat bar, oats, octopus, office chair, olives, onion, onions, orange, organizer, oven, oven glove, paint, paint brush, paint roller, painting, pair, palette knife, pallet, pan, panini press, pants, paper rolls, paper streamers, paper towel, paper towels, paper wrap, paprika, parking sign, parking ticket machine, parsley, party hat, pasta, pea, pen, pencil, pepper, pepper caster, perfume, person, phone, pick, pickaxe, piggy bank, pillow, pills, pine cone, pineapple plant, pineapples, ping pong ball, pipe connection, pipes, pizza, pizza cutter, pizza oven, pizza peel, plane ticket, plant, plastic cups, plate, playing card, plectrum, plug board, plug socket, plunger, pocket balance, pocket knife, pokemon ball, pole, portable electric drill, post, post-its, pot, potato, potato chips, potato masher, potatoes, pouch, pouffe, power drill, powerbank, printer, printer cartridge, projector, puffed rice, pump, pumpkin, puzzle, puzzle piece, rack, racket, radiator, radio, railing, rake, razor, record player, remote, remote control, resistor, ribbon, rice cooker, roll board, roll-on, roller, roller board, rolling pin, rope, rose, rosemary, roses, router, rubber bands, rubber duck, rubik’s cube, running shoes, safe, safety pin, safety pins, safety sign, sage leaves, salad, salami, salmon, salt, sandal, sandpaper, sandwich maker, santa figurine, sausage, sausages, saw, scissor, scissors, scooter, screen, screw, screwdriver, screws, seafood salad, serving tray, serving trolley, sewing machine, shaker, shampoo, shashlik skewers, shaver, shelf, shell, shirt, shoe, shoes, shopping cart, shovel, shovels, shower, shower gel, shower head, shrimp, shuttlecock, sieve, sign, silverware box, sink, skillet, skimmer, skimmers, sledgehammer, sleeping mat, slider, slipper, slippers, sneaker, soap, soap dispenser, sock, socket, socks, spade, speaker, spice mill, spinach, spiral notebooks, sponge, spoon, spout, spray bottle, spray can, spring, squeegee, stamp, stapler, steam flat iron, step bin, stereo, stone plates, stool, stop sign, storage box, straw, straws, string, stuffed animal, suitcase, sunglasses, sunscreen, sweeper, t-shirt, table, takeaway bag, tampon, tap, tape, tape dispenser, tea, tea pot, teabag, tealight, tee box, telephone, tennis ball, terminal, thermal camping mat, thermometer, thread, ticket machine, tin box, tire, tissues, toaster, toilet, toilet bowl, toilet brush, toilet lid, toilet paper, tomato, tomatoes, tongue brush, toothbrush, toothpaste, tortellini, towel, toy car, toy elephant, toy horse, traffic light, traffic signs, trailer, trash bag, trash bin, trash can, trashcan, tray, tree, try square, turmeric, turtle, tv, ufonauts, umbrella, underpants, usb cable, usb-c cable, vacuum, vacuum cleaner, vase, vegetable peeler, video surveillance sign, vinyl, waffles, wall, wall charger, wallet, washbasin, washing machine, watch, water wings, watering can, watermelon, webcam, weight, wet floor sign, wheel, whisk, white cabbage, wildlife camera, window, window cleaner, wine, wood, wooden disc, wooden grill tong, wooden panels, wooden pestle, wooden skewer, wool, wrapping paper, wrench, yarn, and zucchini.
+Table˜10 summarizes the frequency of object labels.
+The top 15 labels (e.g., “lamp”, “plant”, “bottle”) are among the most frequent, while the bottom 5 labels (e.g., “chain”, “roller board”, “mask”) appear rarely.
+Notably, 437 object labels appear exactly once, illustrating a pronounced long-tail distribution that underscores the diversity of the dataset.
+Table 10: Top 15 and bottom 5 object labels by appearance frequency.
+Object label Count
+lamp 13
+plant 13
+bottle 13
+chair 11
+glass 10
+candle 10
+cable 8
+grill 8
+printer 8
+cup 8
+speaker 8
+pillow 8
+brush 7
+canister 7
+pen 7
+chain 1
+roller board 1
+mask 1
+goggles 1
+webcam 1
+SCAM includes a variety of attack words:
+apple, arrow, bag, ball, barrier, basketball, beach, beatles, bed, beer, bell, bench, bicycle, bike, bird, blanket, boat, book, boot, bottle, boulder, bowl, box, boy, bracelet, bridge, bridge out, broom, bus, butter, cable, cake, camera, canal, candle, cane, cannon, cap, car, carpet, carrot, cart, cat, cellphone, chair, cheese, chest, chicken, child, clock, cloud, coat, coin, cone, couch, cow, crab, crystal, cup, curb, curtaxin, curve, dead end, deer, desert, desk, detour, dish, dog, door, duck, eagle, egg, elevator, exit, fan, feather, fence, finger, fire, fish, flag, floor, flower, fog, forest, fork, frog, fruit, garbage bin, gears, ginger, gloves, grandpa, grape, grass, gravel, green, grenade, gun, hair, hammer, hat, helmet, highway, hill, hole, honey, horn, horse, house, ice, iron, jacket, jar, jeep, key, knife, ladder, lake, lamp, lane, leaf, lemon, lentils, lift, lightbulb, lion, lipstick, lock, macbook, mad, median, merge, mirror, mittens, mountaxin, no left turn, no parking, ocean, olive oil, oregano, pedestrian, pen, person, phone, piano, picasso, pig, police, pumpkin, raccoon, red light, risotto rice, river, road, road closed, rock, saddle, salt, sandwich, santa claus, scarf, scooter, screw, screwdriver, sea, seed, shadow, ship, sidewalk, signal, slippery, sloth, slow, smartphone, sofa, speed camera, speed limit, spok, spoon, staxirs, stop, stop sign, store, street, street sign, suv, table, taxi, toll, traffic light, traffic sign, train, transistor, trash, tree, truck, unicorn, van, vase, wall, wet, woman, yield, and zebra.
+Table˜11 summarizes the frequency of attack words.
+The top 15 attack words (e.g., “knife”, “truck”, “slow”) occur frequently, whereas many attack words are rare.
+In fact, 56 attack words appear exactly once, again highlighting overall diversity of the dataset.
+Table 11: Top 15 and bottom 5 attack words by appearance frequency.
+Attack Word Count
+knife 24
+truck 23
+slow 20
+lion 20
+zebra 19
+helmet 18
+bike 17
+honey 15
+pig 15
+camera 15
+lemon 15
+pen 15
+tree 14
+house 14
+fish 13
+beatles 1
+green 1
+raccoon 1
+beach 1
+unicorn 1

### v7

@@
 Finally, we publicly release the datasets introduced in this paper, along with the code for evaluations under www.bliss.berlin/research/scam.
+Keywords: Typographic Attack Dataset, Vision-Language Models, Benchmarking
 1 Introduction
-Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9].
-However, recent research has revealed that these models are vulnerable to typographic attacks [10, 11, 12], which exploit the reliance on the textual component of multimodal learning.
+Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) have demonstrated remarkable performance across various multimodal tasks (Gan et al., 2022), enabling applications in image classification (Radford et al., 2021; Menon and Vondrick, 2022), multimodal retrieval (Bogolin et al., 2022; Ming and Li, 2024), and generative tasks (Ramesh et al., 2022; Rombach et al., 2022; Gafni et al., 2022).
+However, recent research has revealed that these models are vulnerable to typographic attacks (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024), which exploit the reliance on the textual component of multimodal learning.
 Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions.
@@
 Notably, synthetic attacks are as effective as real ones, validating their use for scalable robustness evaluation.
-With the growing use of VLMs in real-world applications, such as medical applications [13, 14, 15] and autonomous systems [16], as well as in Large Vision-Language Models (LVLMs) [17, 18, 19], it is essential to ensure the safety and reliability of these models.
+With the growing use of VLMs in real-world applications, such as medical applications (Zhao et al., 2023; Li et al., 2023b; Hartsock and Rasool, 2024) and autonomous systems (Zhou et al., 2024), as well as in Large Vision-Language Models (LVLMs) (Liu et al., 2024c, a; Dubey et al., 2024), it is essential to ensure the safety and reliability of these models.
 To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date.
@@
   3. 3.
-    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses inherited from their vision encoders.
+    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses present in their vision encoders.
   4. 4.
     Employing larger Large Language Model (LLM) backbones reduces susceptibility to typographic attacks in LVLMs while also improving their typographic understanding.
-Table 1: Performance of VLMs and LVLMs available through OpenCLIP [20] resp.
-ollama [21], and Anthropic’s resp.
+Table 1: Performance of VLMs and LVLMs available through OpenCLIP (OpenCLIP, ) resp.
+ollama (Ollama, ), and Anthropic’s resp.
 OpenAI’s API on the SCAM datasets.
 Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding.
-Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [22].
+Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model (Razzhigaev et al., 2023).
 This table presents a only selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
@@
 Multimodal Foundation Models.
-Multimodal pretraining, as established by CLIP [1] and SigLIP [2], leverages image-text pairs to capture rich semantics suitable for a variety of applications [23, 24].
+Multimodal pretraining, as established by CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), leverages image-text pairs to capture rich semantics suitable for a variety of applications (Dorbala et al., 2022; Zhou et al., 2023).
 This approach commonly features a contrastive objective to align visual and textual content in a shared latent space.
-Following this paradigm, various Vision-Language Models (VLMs) [25, 26, 27] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7].
-Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [28, 29, 30, 31, 32, 33, 34, 35].
-Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
+Following this paradigm, various Vision-Language Models (VLMs) (Jia et al., 2021; Li et al., 2022; Tschannen et al., 2025) have been proposed, demonstrating impressive performance in numerous downstream tasks (Bogolin et al., 2022; Ramesh et al., 2022; Ming and Li, 2024).
+Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks (Hurst et al., 2024; Liu et al., 2023b, a, 2024b; Team et al., 2025; Grattafiori et al., 2024; Meta, 2025; Anthropic, 2025).
+Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation (Rombach et al., 2022; Gafni et al., 2022).
 Typographic Attacks.
-While vision-language pretraining enables strong performance, research has shown that VLMs and LVLMs are susceptible to typographic attacks [10, 11, 12].
-Existing works demonstrate several approaches to employ typographic attacks, for instance by Visual-Prompt-Injection [36], with self-generated attacks [12], or in scene-coherent fashion [37].
-Since VLMs are integrated into autonomous systems, typographic attacks are also known to occur in safety-critical applications such as autonomous driving [38].
+While vision-language pretraining enables strong performance, research has shown that VLMs and LVLMs are susceptible to typographic attacks (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024).
+Existing works demonstrate several approaches to employ typographic attacks, for instance by Visual-Prompt-Injection (Kimura et al., 2024), with self-generated attacks (Qraitem et al., 2024), or in scene-coherent fashion (Cao et al., 2024b).
+Since VLMs are integrated into autonomous systems, typographic attacks are also known to occur in safety-critical applications such as autonomous driving (Chung et al., 2024).
 Datasets.
-To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works [39, 40, 11, 12, 36, 41] synthetically add attack words into existing image datasets while only few [42, 43, 44] propose datasets that exhibit real-world typographic attacks.
-Existing real-world datasets, such as PAINT [42] and Materzynska+[43], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models.
-The previously largest real-world typographic attack dataset, RTA-100 [44], consists of 1000 images spanning 100 objects and attack words.
-However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison).
+To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works (Gong et al., 2023; Cheng et al., 2024d, a, c; Qraitem et al., 2024; Kimura et al., 2024) synthetically add attack words into existing image datasets while only few (Ilharco et al., 2022; Materzyńska et al., 2022; Azuma and Matsui, 2023) propose datasets that exhibit real-world typographic attacks.
+Existing real-world datasets, such as PAINT (Ilharco et al., 2022) and Materzynska+(Materzyńska et al., 2022), lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models.
+The previously largest real-world typographic attack dataset, RTA-100 (Azuma and Matsui, 2023), consists of 1000 images spanning 100 objects and attack words.
+However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison, and section˜A.2 for example images of PAINT and RTA-100).
 Our dataset addresses this limitation by introducing substantially larger variability across object–word combinations, which mitigates dataset bias and enables a more reliable assessment of generalization in VLMs and LVLMs.
@@
 Choosing a uniform font style in this synthetic variant was a deliberate simplification.
-Prior synthetic attack datasets have already explored variations in font position, opacity, and color [40], whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
+Prior synthetic attack datasets have already explored variations in font position, opacity, and color (Cheng et al., 2024a), whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
 Data Annotation and Processing.
 All images were manually annotated with object and attack word labels as well as post-it size (in %) as part of a voluntary community effort.
-The annotation process was conducted using Label Studio [45].
-To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA [46], an interpolation method that preserves details while minimizing distortion.
+The annotation process was conducted using Label Studio (Studio, ).
+To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA (OpenCV, ), an interpolation method that preserves details while minimizing distortion.
 Dataset Statistics.
 SCAM comprises 1162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations.
-The complete lists of objects and attack words are provided in section˜A.8.
+The complete lists of objects and attack words are provided in section˜A.9.
 We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems.
@@
 Dataset Images Objects Attacks
-PAINT [42] 110 89 87
-Materzynska+ [43] 180 20 8
-RTA-100 [44] 1000 100 100
+PAINT (Ilharco et al., 2022) 110 89 87
+Materzynska+ (Materzyńska et al., 2022) 180 20 8
+RTA-100 (Azuma and Matsui, 2023) 1000 100 100
 SCAM (ours) 1162 660 206
 As summarized in LABEL:tab:dataset_comparison, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words.
-Compared to RTA-100 [44], SCAM covers a much broader range of objects (660 vs.
+Compared to RTA-100 (Azuma and Matsui, 2023), SCAM covers a much broader range of objects (660 vs.
 100) and attack words (206 vs.
@@
 The three explicitly aligned components – SCAM, NoSCAM, and SynthSCAM – allow precise, controlled measurement of typographic attack effects and direct comparisons between clean, real, and synthetic conditions, thereby closing critical gaps in prior datasets.
+Maintenance and Ethical Use.
+SCAM is publicly hosted on Hugging Face under a stable organization account [^2] https://huggingface.co/datasets/BLISS-e-V/SCAM.
+The dataset card provides versioning, update logs, and contact information, and we will maintain SCAM by releasing documented updates if errors are discovered or extensions are added.
+Users can contact the corresponding authors to report issues, request corrections, or raise takedown concerns.
+All images in SCAM depict everyday objects with handwritten English words on post-it notes.
+SCAM does not contain personal, biometric, or otherwise sensitive data.
+The dataset is released for research purposes under the license specified on the dataset card, and users are expected to follow standard citation practices and refrain from deploying models trained on SCAM in applications intended to harm or deceive humans.
+A more detailed discussion of potential risks and benefits is provided in the Broader Impact Statement on page Broader Impact Statement.
 4 Evaluation
 We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup.
-Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT [42] and RTA-100 [44].
+Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT (Ilharco et al., 2022) and RTA-100 (Azuma and Matsui, 2023).
 The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
 Experimental Setup.
-To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [47, 48, 1, 49, 20].
-We also evaluate eleven LVLMs, including Claude Sonnet 4 [35], GPT-4o [28], LLaVA [29, 30, 31], Gemma3 [32], Llama3.2-vision [33] and Llama4 [34] models.
-The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform [21], whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp.
+To evaluate VLMs such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), we use the OpenCLIP suite (Ilharco et al., 2021; Schuhmann et al., 2022; Radford et al., 2021; Cherti et al., 2023; OpenCLIP, ).
+We also evaluate eleven LVLMs, including Claude Sonnet 4 (Anthropic, 2025), GPT-4o (Hurst et al., 2024), LLaVA (Liu et al., 2023b, a, 2024b), Gemma3 (Team et al., 2025), Llama3.2-vision (Grattafiori et al., 2024) and Llama4 (Meta, 2025) models.
+The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform (Ollama, ), whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp.
 Anthropic’s API.
@@
 For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word.
-These text embeddings are generated using the set of text templates proposed in [1], as detailed in section˜A.2.
+These text embeddings are generated using the set of text templates proposed in (Radford et al., 2021), as detailed in section˜A.3.
 The predicted label is then determined based on the highest cosine similarity score.
@@
 To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications.
-Our evaluation follows the setup illustrated in fig.˜1c, inspired by [50].
+Our evaluation follows the setup illustrated in fig.˜1c, inspired by (Cheng et al., 2024b).
 The model is provided with an image along with the following prompt:
@@
 […]
-All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and llama3.2-vision:90b and llama4:scout on an H100 80GB.
+All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and
+llama3.2-vision:90b and llama4:scout on an H100 80GB.
+Table 3: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks.
+The tables present average confusion matrices.
+The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks.
+Values represent the mean $\pm$ standard deviation, rounded to integers.
+((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
+SynthSCAM SynthSCAM
+fails succeeds
+SCAM fails $648\pm 192$ $154\pm 68$
+SCAM succeeds $18\pm 11$ $341\pm 151$
+((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
+SynthSCAM SynthSCAM
+fails succeeds
+SCAM fails $741\pm 243$ $96\pm 47$
+SCAM succeeds $124\pm 71$ $201\pm 137$
 5 Results
@@
 LVLMs exhibit similar behavior, with the choice of vision encoder appearing to influence the extent of accuracy degradation.
-Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜4.
+Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜3.
 These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
 datacomp_xl 68.39
-Table 4: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks.
-The tables present average confusion matrices.
-The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks.
-Values represent the mean $\pm$ standard deviation, rounded to integers.
-((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
-SynthSCAM SynthSCAM
-fails succeeds
-SCAM fails $648\pm 192$ $154\pm 68$
-SCAM succeeds $18\pm 11$ $341\pm 151$
-((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
-SynthSCAM SynthSCAM
-fails succeeds
-SCAM fails $741\pm 243$ $96\pm 47$
-SCAM succeeds $124\pm 71$ $201\pm 137$
 Zero-Shot Results for VLMs.
@@
 We investigate the vulnerability of VLMs by examining the impact of their training data, training paradigm, and model architecture, on their susceptibility to typographic attacks.
+Table˜4 presents the impact of training data on the ViT-B models (Dosovitskiy et al., 2020) trained by the OpenCLIP team.
+We focus solely on these models as they share the same training paradigm but are trained on different datasets.
+Overall, the ViT-B models trained on the LAION datasets (Schuhmann et al., 2022) are highly susceptible to typographic attacks, while certain CommonPool variants (Gadre et al., 2023) exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively.
+Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 (Radford et al., 2021) model from OpenAI.
+Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
+While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance.
+We compare different architectures: CLIP-RN (ResNets (He et al., 2016)) and CLIP-ViT models trained by OpenAI (Radford et al., 2021), as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB (Visheratin, 2023) (see table˜5(a)).
+Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models.
+Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
+Moreover, our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table˜5(b).
+Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
 Table 5: VLMs’ architecture impacts susceptibility to typographic attacks.
@@
 The tables present (average) accuracy (Acc.) % on SCAM.
-((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin [51] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1).
-More details about the selection process can be found in section˜A.4.
+((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin (2023) on LAION-COCO NLLB (referred to, in OpenCLIP, as v1).
+More details about the selection process can be found in section˜A.5.
 Training data Architecture Acc. %
@@
 ViT-g-14 14 61.93
-Table˜3 presents the impact of training data on the ViT-B models [52] trained by the OpenCLIP team.
-We focus solely on these models as they share the same training paradigm but are trained on different datasets.
-Overall, the ViT-B models trained on the LAION datasets [47] are highly susceptible to typographic attacks, while certain CommonPool variants [53] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively.
-Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI.
-Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
-While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance.
-We compare different architectures: CLIP-RN (ResNets [54]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [51] (see table˜5(a)).
-Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models.
-Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
-Moreover, our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table˜5(b).
-Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
 Lastly, fig.˜6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks.
-However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
+However, we find that the original CLIP models (Radford et al., 2021) trained by OpenAI perform significantly worse across all sizes.
 Impact of the Size of the Attack.
 We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.˜6).
-The mean accuracy across all models follows this trend, which is consistent with the findings in [40].
+The mean accuracy across all models follows this trend, which is consistent with the findings in (Cheng et al., 2024a).
 Figure 5: Susceptibility to typographic attack (accuracy on SCAM) is agnostic of VLM size, measured in millions of parameters.
@@
 Impact of VLM and LLM Size on LVLM Performance.
-Section˜A.6 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness.
+Section˜A.7 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness.
 The experiment does not show a clear gain in robustness when substituting the vision encoder with one ranked safer by SCAM, while keeping the LLM, training data, and training regime unchanged.
 Instead, the overall performance decreases, highlighting the difficulty of choosing the optimal vision encoder and showcasing the need for further experiments investigating the robustness impact of different VLM choices.
-Apart from that, as discussed in section˜A.7, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.
+Apart from that, as discussed in section˜A.8, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.
 6 Discussion
-Our evaluations on the Subtle Character Attacks on Multimodal Models (SCAM) datasets confirm previous findings [10, 11, 12] that multimodal foundation models remain vulnerable to typographic attacks, whether real-world (SCAM) or synthetic (SynthSCAM).
+Our evaluations on the Subtle Character Attacks on Multimodal Models (SCAM) datasets confirm previous findings (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024) that multimodal foundation models remain vulnerable to typographic attacks, whether real-world (SCAM) or synthetic (SynthSCAM).
 Notably, we find that synthetic attacks are just as effective as real-world ones, leading to comparable performance drops.
@@
 Limitations.
-Despite incorporating a different, “safer” prompt in section˜A.5, a central limitation of our evaluation remains the reliance on “naive prompting” [55, 56] for assessing LVLMs.
-Adopting more advanced multi-prompt evaluation strategies may yield a more reliable measure of model robustness against typographic attacks.
-A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility.
+Despite incorporating a different, “safer” prompt in section˜A.6, a central limitation of our evaluation remains the reliance on “naive prompting” (He et al., 2024; Cao et al., 2024a) for assessing LVLMs.
+More advanced multi-prompt evaluation strategies may yield a more reliable measure of robustness against typographic attacks.
+A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility..
 Future Work.
-As multimodal models are increasingly deployed to safety-critical applications [15, 14, 16, 16], ensuring robustness against typographic attacks is essential to prevent serious consequences.
-Approaches such as alignment training [57] or (mechanistic) defense techniques [44, 58] offer promising directions for enhancing robustness.
-Our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder.
-However, our initial experiments (section˜A.6) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively.
+As multimodal models are increasingly deployed to safety-critical applications (Li et al., 2023b; Hartsock and Rasool, 2024; Zhou et al., 2024), ensuring robustness against typographic attacks is essential to prevent serious consequences.
+Approaches such as alignment training (Afzali et al., 2024) or (mechanistic) defense techniques (Azuma and Matsui, 2023; Hufe et al., 2025) offer promising directions for enhancing robustness.
+The aligned NoSCAM/SCAM pairs are well-suited for causal and mechanistic analyses, as they allow for controlled interventions while keeping scene semantics fixed.
+We view SCAM primarily as a dataset and benchmark to support such work and expect future research to leverage this alignment to develop and evaluate mechanistic defenses against typographic attacks.
+Furthermore, our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder.
+However, our initial experiments (section˜A.7) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively.
 Future work will extend this by systematically retraining LLaVA models with a range of vision backbones to evaluate their role in attack susceptibility and to assess whether encoder strength alone suffices or if additional adaptation steps are necessary.
@@
 We expect that our broad analysis of Vision-Language Models (VLMs) and Large Vision-Language Models (LVLMs) will help researchers make more informed choices when selecting typographically robust models for their safety-critical downstream applications.
+Broader Impact Statement
+Our work highlights a concrete vulnerability in (Large) Vision-Language Models: typographic attacks that cause models to overly rely on embedded text rather than visual content.
+This vulnerability has clear negative implications.
+Malicious actors could exploit our findings to systematically uncover weaknesses in deployed vision–language systems, using typographic cues to mislead models in safety-critical settings such as medical decision support, navigation, or content moderation.
+By releasing a large and diverse dataset of such attacks, we also create the possibility that others could study these loopholes with the explicit aim of designing more effective attacks.
+At the same time, we believe the benefits for research and defense outweigh these risks.
+Robustness research requires shared realistic benchmarks and reproducible baselines.
+SCAM provides a standardized resource that allows the community to evaluate typographic vulnerabilities and compare defense methods systematically, and to study how training data and model design affect susceptibility.
+By demonstrating that synthetic attacks reliably mirror real-world ones, we also reduce the need for ongoing collection of handwritten adversarial examples, lowering barriers for future research.
 Acknowledgments and Disclosure of Funding
@@
 Figure 7: Example images of SCAM, NoSCAM, and SynthSCAM.
-A.2 Zero-Shot Evaluation of LVMs: Text Templates
+A.2 Example Images of PAINT and RTA-100
+Although qualitative inspection is necessarily subjective and 20 samples from the PAINT (Ilharco et al., 2022) resp.
+RTA-100 (Azuma and Matsui, 2023) datasets in figs.˜9 and 9 cannot capture their full variability, the quantitative evidence in LABEL:tab:dataset_comparison and the attack-word distribution shown in fig.˜3 clarify the structural differences.
+PAINT and RTA-100 cover comparatively restricted sets of objects, attack words, and scene conditions.
+SCAM, by contrast, comprises 1162 images with 660 objects and 206 attack words collected by nine contributors across diverse indoor and outdoor environments.
+The resulting variation in lighting, backgrounds, camera quality, handwriting styles, and viewpoints is visible in fig.˜2, even if such diversity is difficult to discern from limited qualitative examples alone.
+Figure 8: 20 random example images from the PAINT (Ilharco et al., 2022) dataset. Best viewed zoomed in.
+Figure 9: 20 random example images from the RTA-100 (Azuma and Matsui, 2023) dataset. Best viewed zoomed in.
+A.3 Zero-Shot Evaluation of LVMs: Text Templates
 To compute the cosine similarity between an image and text embeddings in our zero-shot evaluation of Vision-Language Models (VLMs), we generate text embeddings for both the object label and the attack word using multiple textual variations.
-Specifically, we apply 32 diverse text templates as in [1], each incorporating the object label resp.
+Specifically, we apply 32 diverse text templates as in (Radford et al., 2021), each incorporating the object label resp.
 attack word in a different context.
@@
 The templates are as follows: ( ’a bad photo of a {}.’, ’a photo of many {}.’, ’a sculpture of a {}.’, ’a photo of the hard to see {}.’, ’a low resolution photo of the {}.’, ’a rendering of a {}.’, ’graffiti of a {}.’, ’a bad photo of the {}.’, ’a cropped photo of the {}.’, ’a tattoo of a {}.’, ’the embroidered {}.’, ’a photo of a hard to see {}.’, ’a bright photo of a {}.’, ’a photo of a clean {}.’, ’a photo of a dirty {}.’, ’a dark photo of the {}.’, ’a drawing of a {}.’, ’a photo of my {}.’, ’the plastic {}.’, ’a photo of the cool {}.’, ’a close-up photo of a {}.’, ’a black and white photo of the {}.’, ’a painting of the {}.’, ’a painting of a {}.’, ’a pixelated photo of the {}.’, ’a sculpture of the {}.’, ’a bright photo of the {}.’, ’a cropped photo of a {}.’, ’a plastic {}.’, ’a photo of the dirty {}.’, ’a jpeg corrupted photo of a {}.’, ’a blurry photo of the {}.’ ).
-A.3 Model Accuracies on Typographic Attack Datasets
-Table 6: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on SCAM, PAINT [42], and RTA-100 [44] datasets.
+A.4 Model Accuracies on Typographic Attack Datasets
+Table 6: Performance of (Large) Vision-Language Models available through OpenCLIP (OpenCLIP, ) resp.
+ollama (Ollama, ), and Anthropic’s resp.
+OpenAI’s API on SCAM, PAINT (Ilharco et al., 2022), and RTA-100 (Azuma and Matsui, 2023) datasets.
 Model Training data Accuracy (%)
@@
 gpt-4o-2024-08-06 - 99.48 96.82 94.58 99.09 96.73
-A.4 Architecture Comparison
+A.5 Architecture Comparison
 In table˜5(a) in the main text, we compare different architectures (CLIP-RN vs.
@@
 The table groups models by architecture category and indicates the corresponding fixed training data.
-In OpenCLIP, v1 refers to LAION-COCO NLLB [51].
+In OpenCLIP, v1 refers to LAION-COCO NLLB (Visheratin, 2023).
 Architecture Category Model Training Data
@@
 nllb-clip-large v1
-A.5 Prompt Variation
-Our evaluation of LVLMs in section˜4 focuses on a single prompt.
-Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation [59, 50], we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt.
+A.6 Prompt Variation
+Our evaluation of Large Vision-Language Models (LVLMs) in section˜4 focuses on a single prompt.
+Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation (Mizrahi et al., 2024; Cheng et al., 2024b), we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt.
 Contrary to the expectation formulated in these works, this modification does not reduce vulnerability.
-As illustrated in fig.˜8, the susceptibility to typographic attacks remains largely unchanged.
+As illustrated in fig.˜10, the susceptibility to typographic attacks remains largely unchanged.
 Figure 10: The modified “safer” prompt does not alleviate susceptibility to typographic attacks.
-A.6 Replacing VLM in LLaVA
-Previous work [40, 60] suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact.
-Further, [61] demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in [60].
-Our evaluations on SCAM support the conclusion that some VLMs exhibit higher typographic robustness than others.
-In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA [14, 30, 31] – show heightened vulnerability.
-To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder [62, 63], which was trained on datacomp1b and differs primarily in training data and regime, not architecture.
+A.7 Replacing VLM in LLaVA
+Previous work (Cheng et al., 2024a; Li et al., 2024) suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact.
+Further, (Malik et al., 2025) demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in (Li et al., 2024).
+Our evaluations on SCAM support the conclusion that some Vision-Language Models (VLMs) exhibit higher typographic robustness than others.
+In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA (Li et al., 2023b; Liu et al., 2023a, 2024b) – show heightened vulnerability.
+To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder (X.
+Li, Z.
+Wang, and C.
+Xie (2023c); 13), which was trained on datacomp1b and differs primarily in training data and regime, not architecture.
 Compared to the OpenAI encoder, the CLIPA model exhibits a substantially lower robustness drop: only 25% from NoSCAM to SCAM, versus 65% (see table˜1).
-It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP [48] [^3] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
+It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP (Ilharco et al., 2021) [^4] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
 […]
 Table˜8 shows that our CLIPA-based LLaVA variant does not outperform the original model on SCAM, despite using a VLM that is more robust in isolation.
-Performance on other benchmarks evaluated via VLMEvalKit [65] also drops.
+Performance on other benchmarks evaluated via VLMEvalKit (Duan et al., 2024) also drops.
 These findings suggest that the benefits of a more robust vision encoder do not transfer under naive LLaVA integration.
@@
 Despite the CLIPA vision encoder showing stronger robustness in isolation, this does not translate to improved performance on SCAM.
-Evaluations for the upper section are conducted using VLMEvalKit [65].
+Evaluations for the upper section are conducted using VLMEvalKit (Duan et al., 2024).
 Other results follow the protocol outlined in the main part of the paper.
-Benchmark LLaVA-v1.5-7B (original) LLaVA-v1.5-7B-CLIPA (ours)
-MMBench v1.1 [66] 64.32 60.22
-MME Perception [67] 1444.81 1378.96
-SEED-Bench-1 (img) [68] 66.81 64.02
-TextVQA [69] 46.93 45.57
-OCRBench [70] 31.90 31.60
-OCR-VQA [71] 60.64 57.39
+Benchmark LLaVA-v1.5-7B (orig.) LLaVA-v1.5-7B-CLIPA (ours)
+MMBench v1.1 (Liuand et al., 2023) 64.32 60.22
+MME Perception (Fu et al., 2023) 1444.81 1378.96
+SEED-Bench-1 (img) (Li et al., 2023a) 66.81 64.02
+TextVQA (Singh et al., 2019) 46.93 45.57
+OCRBench (Liu et al., 2024d) 31.90 31.60
+OCR-VQA (Mishra et al., 2019) 60.64 57.39
 NoSCAM 97.16 89.66
@@
 SynthSCAM 51.55 46.08
-PAINT [42] 52.73 53.64
-RTA100 [44] 48.30 47.30
-A.7 Typographic Robustness vs. Typographic Understanding
+PAINT (Ilharco et al., 2022) 52.73 53.64
+RTA100 (Azuma and Matsui, 2023) 48.30 47.30
+A.8 Typographic Robustness vs. Typographic Understanding
 The main paper shows that increasing the size of LLM backbones reduces susceptibility to typographic attacks (see results for LLaVA and Gemma3 in table˜1).
@@
 Together, these findings suggest that typographic robustness emerges as model size increases, alongside improved textual perception.
-However, [36] reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
-Table 9: Performance of selected LLaVA-Next [31] models on three OCR-related benchmarks (TextVQA [69], OCRBench [70] (from VLMEvalKit [65]), and OCR-VQA [71]) shows that increasing LLM backbone size leads to improved typographic understanding.
+However, (Kimura et al., 2024) reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
+Table 9: Performance of selected LLaVA-Next (Liu et al., 2024b) models on three OCR-related benchmarks (TextVQA (Singh et al., 2019), OCRBench (Liu et al., 2024d) (from VLMEvalKit (Duan et al., 2024)), and OCR-VQA (Mishra et al., 2019)) shows that increasing LLM backbone size leads to improved typographic understanding.
 All models use the ViT-L-14-336 vision encoder.
-Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit [65].
+Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit (Duan et al., 2024).
 Model Params (B) TextVQA (%) OCRBench OCR-VQA (%)
@@
 LLaVA-Next-Vicuna-7B 7.1 63.8 532 63.8
-A.8 SCAM Labels
+A.9 SCAM Labels
 The SCAM dataset includes diverse object labels: